### v1

# bibliography.bib

Abstract.
In Outline of a Theory of Truth, Kripke introduces some of the central concepts of the logical study of truth and paradox. He informally defines some of these—such as groundedness and paradoxicality—using modal locutions. We introduce a modal language for regimenting these informal definitions. Though groundedness and paradoxicality are expressible in the modal language, we prove that intrinsicality—which Kripke emphasizes but does not define modally—is not. We characterize the modally definable relations and completely axiomatize the modal semantics.
Thanks to Dan Appel, Reid Dale, Noah Schweber, Ted Slaman, and Kentarô Yamamoto for helpful discussion. Thanks to Neil Barton, Wes Holliday, and Andrew Lee for comments on drafts.
1. Introduction
In Outline of a Theory of Truth, Kripke writes that
It has long been recognized that some of the intuitive trouble with Liar sentences is shared with such sentences as [the truth-teller, i.e., “This sentence is true”] which, though not paradoxical, yield no determinate truth conditions. ([kripke1975outline], p. 693)
So paradoxicality is just one of the troubling features that self-referential sentences can exhibit. Others are ungroundedness and non-intrinsicality. One of the great successes of Kripke’s theory is that it facilitates a classification of sentences according to these properties. Kripke informally describes groundedness and paradoxicality using modal locutions.
…[The truth-teller] is ungrounded, but not paradoxical. This means that we could consistently use the predicate ‘true’ so as to give [the truth-teller] a truth value, though [we need] not do so.…[The liar] cannot have a truth value…. We could consistently use the word ‘true’ so as to give a truth value to such a sentence as [the truth-teller] …. The same does not hold for the paradoxical sentences. ([kripke1975outline], p. 708, emphasis added [^1] Except to the first instance of ‘could’, which is emphasized in the original.)
Yet Kripke does not describe the intrinsic sentences using modal locutions. This is a striking contrast. Is it a mere coincidence that modal locutions appear in the description of the grounded and paradoxical sentences but not of the intrinsic sentences? Or does it reflect some inherent discrepancy between the objects being defined?
The standard way of studying these objects obviates this question. In particular, Kripke replaces the informal modal glosses on paradoxicality and groundedness with official extensional mathematical definitions. Nevertheless, the modal glosses are nice examples of modality occurring naturally in a logico-mathematical context, so it is worth regimenting them in a way that retains their modal flavor. For the purposes of taking these modal definitions “at face value,” we introduce a modal language and an attendant interpretation $\mathbf{FIX}$. Our modal language is propositional, but it contains structured atomic formulas (like in first-order logic) as opposed to propositional variables. In particular, there are atomic formulas of the form:
$T(x)\quad\text{for ``$x$ is true''}$”
$F(x)\quad\text{for ``$x$ is false''.}$”.
We also use $N(x)$ is an abbreviation for $\neg T(x)\wedge\neg F(x)$. Kripke’s informal modal definitions of groundedness and paradoxicality are straightforwardly formalizable in this setting.
We provide a complete and decidable axiomatization of the modal semantics. We introduce the following axioms:
$\mathbf{Con}$$\quad\neg\big{(}T(x)\wedge F(x)\big{)}$
$\mathbf{Ground}$ $\quad(\lozenge T(x)\wedge\lozenge F(x))\to\lozenge N(x)$
$\mathbf{Min_{n}}$ $\quad\big{(}\lozenge N(x_{1})\wedge\dots\wedge\lozenge N(x_{n})\big{)}\to\lozenge\big{(}N(x_{1})\wedge\dots\wedge N(x_{n})\big{)}$
and define $\mathbf{S5[Con,Ground,Min_{n}]}$ as the normal modal logic that results from adjoining these axioms to $\mathbf{S5}$. Our main completeness theorem is the following:
Theorem.
For all formulas $\varphi$ with $n$ variables, the following are equivalent:
  1. (i)
    $\mathbf{FIX}\Vdash\varphi$
  2. (ii)
    $\mathbf{S5[Con,Ground,Min_{n}]}\vdash\varphi$
This completeness theorem falls out of an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$. From this characterization, we infer that the set of intrinsic sentences is not modally indefinable. Thus, it is not a mere coincidence that Kripke does not lapse into modal language when defining intrinsicality. Indeed, we are able to show that the intrinsic sentences are indiscernible within the modal language from another set, namely, the set of inevitable sentences, which we define later in the paper. The characterization of the modally definable sets builds on a classic normal form theorem for $\mathbf{S5}$ from [carnap1946modalities]. According to Carnap’s theorem, every formula is $\mathbf{S5}$-equivalent to a disjunction of formulas in a very special form. This normal form result and its corollaries adumbrated here resemble classic results in model theory (particular, results in o-minimality and quantifier elimination). We conclude with a discussion of this analogy.
2. Kripke’s Theory of Truth
[…]
Let $\mathcal{L}$ be the first-order language of arithmetic. Let $\mathcal{L}_{T}$ be the language of arithmetic augmented with the unary predicate $\mathsf{True}(x)$. Some presentations also include a predicate $\mathsf{False}(x)$, but one may interpret falsity as truth-of-negation, whence the second predicate is strictly unnecessary. As usual, $\ulcorner\varphi\urcorner$ is the Gödel number of the $\mathcal{L}_{T}$ sentence $\varphi$ according to some standard Gödel numbering.
2.1. Strong Kleene logic
Kripke defines his interpretations of $\mathcal{L}_{T}$ using the trivalent logic $\mathsf{K3}$, also known as the strong Kleene logic. According to $\mathsf{K3}$, a structure $w$ can stand in one of three semantic relations with a formula $\varphi$:
  • $w\vDash_{\mathsf{K3}}\varphi$, i.e., $\varphi$ is true in $w$
  • $w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}\varphi$, i.e., $\varphi$ is false in $w$
  • $w\uparrow_{\mathsf{K3}}\varphi$ i.e., $\varphi$ is neither in $w$
The semantics of conjunction is given by the following clauses:
  • $w\vDash_{\mathsf{K3}}\varphi\wedge\psi$ if $w\vDash_{\mathsf{K3}}\varphi$ and $w\vDash_{\mathsf{K3}}\varphi$
  • $w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}\varphi\wedge\psi$ if $w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}\varphi$ or $w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}\psi$
  • $w\uparrow_{\mathsf{K3}}\varphi$ otherwise.
The semantics of negation is given by the following clauses:
  • $w\vDash_{\mathsf{K3}}\neg\varphi$ if $w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}\varphi$
  • $w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}\neg\varphi$ if $w\vDash_{\mathsf{K3}}\varphi$
  • $w\uparrow_{\mathsf{K3}}\neg\varphi$ otherwise.
The semantics of the universal quantifier is given by the following clauses:
  • $w\vDash_{\mathsf{K3}}\forall x\;\varphi(x)$ if for every $d\in w$, $w\vDash_{\mathsf{K3}}\varphi(d)$.
  • $w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}\forall x\;\varphi(x)$ if for some $d\in w$, $w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}\varphi(d)$.
  • $w\uparrow_{\mathsf{K3}}\forall x\;\varphi(x)$ otherwise.
2.2. Classification of sentences
Kripke introduces a large collection of 3-valued (true, false, neither) models for $\mathcal{L}_{T}$. These 3-valued models are typically called “fixed points,” since Kripke defines them as the fixed points of a certain function (namely, the Kripke jump). The details of this function are not important for us, but the “fixed points” terminology is ubiquitous, so we will continue to use it.
Remark 2.1.
Each fixed point $w$ has the following properties:
  1. (1)
    For each $\varphi$, $\varphi$ and $\mathsf{True}(\ulcorner\varphi\urcorner)$ have the same semantic value in $w$.
  2. (2)
    $w$ is consistent, i.e., there is no $\varphi$ such that $w\Vdash_{\mathsf{K3}}\varphi$ and $w\Vdash_{\mathsf{K3}}\neg\varphi$.
  3. (3)
    Hence, there is no $\varphi$ such that $w\Vdash_{\mathsf{K3}}\mathsf{True}(\ulcorner\varphi\urcorner)\wedge\mathsf{True}(\ulcorner\neg\varphi\urcorner)$
For any formula $\varphi(x)$ in the language of $\mathcal{L}_{T}$, we can formulate an $\mathcal{L}_{T}$ sentence $\psi$ which “says” $\varphi(\ulcorner\psi\urcorner)$. In particular, $\psi$ and $\varphi(\ulcorner\psi\urcorner)$ have the same semantic value in each fixed point. For instance, there are $\mathcal{L}_{T}$ sentences $\lambda,\tau,\gamma$ such that:
  • $\lambda$ has the same semantic value as $\neg\mathsf{True}(\ulcorner\lambda\urcorner)$ in every fixed point.
  • $\tau$ has the same semantic value as $\mathsf{True}(\ulcorner\tau\urcorner)$ in every fixed point.
  • $\gamma$ has the same semantic value as $\mathsf{True}(\ulcorner\gamma\urcorner)\vee\neg\mathsf{True}(\ulcorner\gamma\urcorner)$ in every fixed point.
In fact, there are many such sentences. We will sometimes appeal to this fact, which is well-known in the theory of truth literature:
Remark 2.2.
There are infinitely many logically independent truth-teller sentences. That is, there is an infinite collection $T$ of sentences such that:
  • For all $\tau\in T$, $\tau$ has the same semantic value as $\mathsf{True}(\ulcorner\tau\urcorner)$ in every fixed point.
  • For any sequence $\tau_{1},\dots,\tau_{n}$ of sentences from $T$ and any assignment of true, false, and neither to $\tau_{1},\dots,\tau_{n}$, there is a fixed point realizing exactly that truth-value assignment.
Definition 2.3.
A sentence is grounded just in case (i) it has the value true in all models or (ii) it has the value false in all models.
Remark 2.4.
There is a minimum fixed point in which all ungrounded sentences have the value neither. Hence, a sentence is ungrounded just in case it has the value neither in some fixed point. The following claims follow:
  1. (1)
    If a sentence has value true in one fixed point and false in another, then it has value neither in the minimum fixed point.
  2. (2)
    If $\varphi_{1},\dots,\varphi_{n}$ are ungrounded, then they all have value neither in the minimum fixed point.
The liar $\lambda$ , truth-teller $\tau$, and tautology-teller $\gamma$ are all ungrounded, so they all have value neither in the minimum model.
Definition 2.5.
A sentence is paradoxical just in case it has value neither in all models.
The liar $\lambda$ is paradoxical. But the truth-teller $\tau$ and tautology-teller $\gamma$ are not paradoxical:
  • $\tau$ has value true in some models and value false in others.
  • $\gamma$ has value true in some models and does not have the value false in any models.
Though neither $\tau$ nor $\gamma$ is paradoxical, it is worth noting that $\gamma$ alone cannot take on both of the binary truth values. We introduce the following definition to discuss this property of $\gamma$:
Definition 2.6.
A sentence is inevitable if it has one of the binary truth-values in some fixed point and the other binary truth-value in no fixed point.
Any grounded sentence is inevitable. There are also ungrounded inevitable sentences. For instance, if $\tau$ is a truth-teller, then the sentence $\tau\vee\neg\tau$ is inevitable (inevitably true) as is $\tau\wedge\neg\tau$ (inevitably false).
Informally, the intrinsic sentences are those inevitable sentences whose truth/falsity does not depend on any arbitrary sentences. Here is a more precise definition:
Definition 2.7.
A fixed point $w$ is intrinsic if for every fixed point $v$,
[…]
is consistent. A sentence $\varphi$ is intrinsic if it is true or false in an intrinsic fixed point.
Note that inevitability and intrinsicality come apart. The tautology-teller $\gamma$ is both intrinsic and inevitable. By contrast, $\tau\vee\neg\tau$ is inevitable but not intrinsic. The issue is that any fixed point which makes $\tau\vee\neg\tau$ true must either make $\tau$ true or $\neg\tau$ true. So $\tau\vee\neg\tau$ can only be made true by making a binary truth-value assignment that is incompatible with other possible binary truth-value assignments.
3. Modal Systems
In this section we introduce our modal framework, including the language, semantics, and axiom systems. Let’s start with a presentation of the modal language. In what follows:
  • $\mathcal{X}$ is a countably infinite set of variables.
  • $T$ and $F$ are one-place predicate symbols.
  • $\neg$, $\wedge$, and $\Box$ are sentential connectives.
Definition 3.1.
The formulas of $\mathcal{L}_{\Box}$ are defined as follows:
$\varphi::=T(x)\mid F(x)\mid\neg\varphi\mid(\varphi\wedge\varphi)\mid\Box\;\varphi$
That is, a formula is a predicate applied to a variable, a negation of a formula, a conjunction of formulas, or a box appended to a formula. Though ,,$\vee$,$\to$,$\leftrightarrow$,$\lozenge$ , → , ↔ , ◆ are not officially part of the language we sometimes use these symbols as abbreviations in the standard way. Note that the language $\mathcal{L}_{\Box}$ neither contains quantifiers nor propositional variables.
For the rest of this paper we will be working with this signature, i.e., with predicates $T$ and $F$. For instance, when we discuss the modal logic $\mathbf{S5}$, we mean the modal logic $\mathbf{S5}$ in this signature.
Definition 3.2.
$\mathbf{S5}$ is the classical modal logic containing all instances of the following axioms:
$\mathbf{K}$ $\quad\Box(A\to B)\to(\Box A\to\Box B)$
$\mathbf{T}$ $\quad\Box A\to A$
$\mathbf{5}$ $\quad\lozenge A\to\Box\lozenge A$
and closed under the following rule:
$\mathbf{RN}$ If A is a theorem, so is $\quad\text{If $A$ is a theorem, so is $\Box A$}.$
We will now introduce models for interpreting the formal system $\mathbf{S5}$.
Definition 3.3.
For a set $W$, a variable assignment $V$ is a pair of functions $V_{1},V_{2}:\mathcal{X}\to\mathcal{P}(W)$.
Definition 3.4.
Given $w\in W$ and variable assignment $V$, we define $\Vdash_{V}$ as follows:
  • $W,w\Vdash_{V}T(x)$ if $w\in V_{1}(x)$.
  • $W,w\Vdash_{V}F(x)$ if $w\in V_{2}(x)$.
  • $W,w\Vdash_{V}\neg\varphi$ if $W,w\nVdash_{V}\varphi$.
  • $W,w\Vdash_{V}\varphi\wedge\psi$ if $W,w\Vdash_{V}\varphi$ and $W,w\Vdash_{V}\psi$.
  • $W,w\Vdash_{V}\Box\;\varphi$ if for all $v\in W$, $W,v\Vdash_{V}\varphi$.
We say that $W\Vdash_{V}\varphi$ if for all $w\in W$, $W,w\Vdash_{V}\varphi$.
We will be particularly interested in a particular $\mathbf{S5}$ structure.
Definition 3.5.
Let $\mathbf{FIX}$ be the set of Kripke’s fixed points, i.e., the set of Kripke’s 3-valued $\mathcal{L}_{T}$ models. An $\mathcal{L}_{T}$-realization is a function $\star:\mathcal{X}\to\mathcal{L}_{T}$, i.e., a mapping from variables to $\mathcal{L}_{T}$ formulas. Each $\mathcal{L}_{T}$-realization $\star$ determines a variable assignment $V^{\star}$:
  1. (i)
    $V^{\star}_{1}(x)$ is the set of models in which $x^{\star}$ is true.
  2. (ii)
    $V^{\star}_{2}(x)$ is the set of models in which $x^{\star}$ is false.
$\mathbf{FIX}\Vdash\varphi$ just in case $\mathbf{FIX}\Vdash_{V^{\star}}\varphi$ for every $\mathcal{L}_{T}$-realization $\star$.
For any $\star$, $\mathbf{FIX}\Vdash_{V^{\star}}\mathbf{S5}$. In fact, $\mathbf{FIX}$ validates other principles that can be stated in the modal language $\mathcal{L}_{\Box}$.
Definition 3.6.
$\mathbf{S5[Con]}$ extends $\mathbf{S5}$ with the following axiom:
$\mathbf{Con}\quad\neg\big{(}T(x)\wedge F(x)\big{)}.$
In what follows, $N(x)$ is an abbreviation for the formula:
$\neg T(x)\wedge\neg F(x).$
Definition 3.7.
$\mathbf{S5[Con,Ground]}$ extends $\mathbf{S5[Con]}$ with the following axiom:
$\mathbf{Ground}\quad(\lozenge T(x)\wedge\lozenge F(x))\to\lozenge N(x).$
We will introduce more axioms as they are needed. For now we note the following.
Proposition 3.8.
For any $\star$, $\mathbf{FIX}\Vdash_{V^{\star}}\mathbf{S5[Con,Ground]}$.
Proof.
By Remark 2.1 and Remark 2.4. ∎
4. Single Variable Definability
The motivations discussed in the introduction concerned definability of sets, such as the set of grounded sentences. To study definability of sets, it suffices to focus on 1-formulas (i.e., formulas with a single variable). Accordingly, in this section we will limit our attention to 1-formulas. We will turn to $n$-formulas for $n>1$ in subsequent sections. Characterizing the definable sets is much easier than characterizing the definable $n$-ary relations, so this section will also serve as a warm up for the later sections.
Definition 4.1.
An $\mathcal{L}_{\Box}$ formula $\varphi(x)$ defines $\mathcal{A}$ over $\mathbf{FIX}$ just in case for every $\mathcal{L}_{T}$-realization $\star$:
$\mathbf{FIX}\Vdash_{V^{\star}}\varphi(x)\Longleftrightarrow x^{\star}\in\mathcal{A}.$
Let’s provide some examples:
$x\text{ is grounded }$ $:=\Box\;T(x)\vee\Box\;F(x)$
$x\text{ is paradoxical }$ :=$:=\neg\lozenge T(x)\wedge\neg\lozenge F(x)$
$x\text{ is inevitable }$ $:=\big{(}\lozenge T(x)\wedge\neg\lozenge F(x)\big{)}\vee\big{(}\neg\lozenge T(x)\wedge\lozenge F(x)\big{)}$
One conspicuous absence from the list is intrinsicality. This is not an oversight. One goal of this section is to show that the intrinsic sentences have no modal definition.
4.1. Normal Forms
One of our main tools will be a normal form theorem due to [carnap1946modalities]. Let’s begin by drawing a distinction between two kinds of formulas.
Definition 4.2.
An extensional formula is one without any modal operators. An intensional formula is one all of whose variables occurs only in the scope of a modal operator.
Note that the distinction between extensional and intensional formulas is exclusive but not exhaustive.
Definition 4.3.
Let $\Gamma$ be a set of formulas and $T$ a modal system. A formula $\varphi$ is $\Gamma$-maximal for $T$ if each of the following holds:
  1. (1)
    $T+\varphi$ is consistent
  2. (2)
    $\varphi\in\Gamma$
  3. (3)
    For every $\psi\in\Gamma$, one of the following holds:
    1. (a)
      $T\vdash\varphi\to\psi$
    2. (b)
      $T\vdash\varphi\to\neg\psi$
The following notion is inspired by the model-theoretic notion of an isolated type:
Definition 4.4.
The extensional 1-isolators for $T$ are the $\Gamma$-maximal formulas for $T$ where $\Gamma$ is the set of extensional 1-formulas.
Proposition 4.5.
The extensional 1-isolators for $\mathbf{S5}$:
  1. (i)
    $T(x)\wedge F(x)$
  2. (ii)
    $T(x)\wedge\neg F(x)$
  3. (iii)
    $\neg T(x)\wedge F(x)$
  4. (iv)
    $\neg T(x)\wedge\neg F(x)$
$\mathbf{S5[Con]}$ refutes (i) and makes the additional conjuncts in (ii) and (iii) redundant. Hence, we have the following:
Proposition 4.6.
The extensional 1-isolators for $\mathbf{S5[Con]}$ are:
$T(x),F(x),N(x)$
There is an analogous notion of intensional 1-isolators:
Definition 4.7.
The intensional 1-isolators for $T$ are the $\Gamma$-maximal formulas for $T$ where $\Gamma$ is the set of intensional 1-formulas.
Let’s restrict our attention to $\mathbf{S5[Con]}$ and its extensions. Here is how we characterize the intensional 1-isolators of such a system. We begin by considering all formulas of the form:
¬${}^{+}_{\neg}\lozenge T(x)\wedge^{+}_{\neg}\lozenge F(x)\wedge^{+}_{\neg}\lozenge N(x)$
That is, for each of the extensional 1-isolators, we choose whether it is possible or not and add that choice as a conjunct. Call these the intensional pre-1-isolators. To isolate the intensional 1-isolators we must figure out which of the intensional pre-1-isolators are consistent.
$\mathbf{S5[Con]}$ refutes exactly one of the intensional pre-1-isolators.
$\mathbf{S5[Con]}\vdash\neg\big{(}\neg\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\neg\lozenge N(x)\big{)}$
$\mathbf{S5[Con,Ground]}$ refutes one more in addition:
$\mathbf{S5[Con,Ground]}\vdash\neg\big{(}\lozenge T(x)\wedge\lozenge F(x)\wedge\neg\lozenge N(x)\big{)}$
Definition 4.8.
A 1-isolator for $T$ is a $T$-consistent conjunction $\varphi\wedge\psi$ where $\varphi$ is an extensional 1-isolator and $\psi$ is an intensional 1-isolator.
We are now ready to state Carnap’s theorem. To make sense of Carnap’s theorem, recall the convention that an empty disjunction of formulas is just a contradiction.
Theorem 4.9 (essentially [carnap1946modalities]).
For $T$ extending $\mathbf{S5}$:
  1. (1)
    Every extensional 1-formula is $T$ equivalent to a disjunction of extensional 1-isolators for $T$.
  2. (2)
    Every intensional 1-formula is $T$ equivalent to a disjunction of intensional 1-isolators for $T$.
  3. (3)
    Every 1-formula is $T$ equivalent to a disjunction of 1-isolators for $T$.
4.2. Counting
[carnap1946modalities] calculates the number of formulas up to provable equivalence in the modal logic $\mathbf{S5}$. Note the emphasis on the word logic. The system that Carnap studies is in the typical language of modal logic, i.e., without the predicates $T,F$ but with propositional variables $p_{1},p_{2},\dots$. There are far more equivalence classes of 1-formulas in our system $\mathbf{S5}$ than in the modal logic Carnap studied.
Proposition 4.10.
Up to $\mathbf{S5}$-equivalence, there are exactly $4{,}294{,}967{,}296$ many 1-formulas.
Proof.
Let’s introduce some abbreviations:
$B(x)$:
$T(x)\wedge F(x)$
$T_{only}(x)$:
$T(x)\wedge\neg F(x)$
$F_{only}(x)$:
$\neg T(x)\wedge F(x)$
$N(x)$:
$\neg T(x)\wedge\neg F(x)$
$\mathbf{S5}$ has 4 extensional 1-isolators and 15 intensional 1-isolators, namely, all formulas of the form:
¬${}^{+}_{\neg}\lozenge B(x)\wedge^{+}_{\neg}\lozenge T_{only}(x)\wedge^{+}_{\neg}\lozenge F_{only}(x)\wedge^{+}_{\neg}\lozenge N(x)$
except for the one with all negations.
Each extensional 1-isolator is inconsistent with the intensional 1-isolators claiming that it is impossible, and consistent otherwise.
So, for instance, $B(x)$ is consistent with all and only those intensional 1-isolators of the form:
$\lozenge B(x)\wedge^{+}_{\neg}\lozenge T_{only}(x)\wedge^{+}_{\neg}\lozenge F_{only}(x)\wedge^{+}_{\neg}\lozenge N(x)$
of which there are exactly $82^{3}=$ 8.
A 1-isolator is a consistent conjunction of an extensional 1-isolator and an intensional 1-isolator. So there are $4\times 8=32$ many 1-isolators for $\mathbf{S5}$. Hence, the number of formulas up to provable equivalence is exactly $2962^{32}=4{,}294{,}967{,}$ 296. ∎
Proposition 4.11.
Up to $\mathbf{S5[Con]}$-equivalence, there are exactly $4096$ 1-formulas.
Proof.
We approach this calculation in a similar way. This time the extensional 1-isolators are $T(x),F(x),N(x)$.
Focus on $T(x)$ for a moment. It is consistent with all and only those intensional 1-isolators of the form:
$\lozenge T(x)\wedge^{+}_{\neg}\lozenge F(x)\wedge^{+}_{\neg}\lozenge N(x)$
of which there are exactly $42^{2}=$ 4.
A 1-isolator is a consistent conjunction of an extensional 1-isolator and an intensional 1-isolator. So there are $3\times 4=12$ many 1-isolators for $\mathbf{S5[Con]}$. Hence, the number of formulas up to provable-equivalence is exactly $40962^{12}=$ 4096. ∎
Proposition 4.12.
Up to $\mathbf{S5[Con,Ground]}$-equivalence, there are exactly $0241$-formulas.
Proof.
Again, the extensional 1-isolators are $T(x),F(x),N(x)$.
Focus on $T(x)$ for a moment. It is consistent with all and only those intensional 1-isolators of the form:
$\lozenge T(x)\wedge^{+}_{\neg}\lozenge F(x)\wedge^{+}_{\neg}\lozenge N(x)$
except (by the $\mathbf{Ground}$ axiom) for
$\lozenge T(x)\wedge\lozenge F(x)\wedge\neg\lozenge N(x).$
So $T(x)$ is consistent with $32^{2}-1=4-1=$ 3 intensional 1-isolators. Likewise, $F(x)$ is consistent with $3$ intensional 1-isolators. $N(x)$ is consistent with $4$ (there is no need to delete one since we focus on formulas including $\lozenge N(x)$).
A 1-isolator is a consistent conjunction of an extensional 1-isolator and an intensional 1-isolator. So there are $3+3+4=10$ many 1-isolators for $\mathbf{S5[Con]}$. Hence, the number of formulas up to provable-equivalence is exactly $10242^{10}=$ 1024. ∎
4.3. Completeness
In this subsection we present a complete axiomatization of the one-variable fragment of the modal theory of $\mathbf{FIX}$. Completeness follows easily from Carnap’s normal form theorem. The connection between normal forms and completeness theorems was explored previously in [fine1975normal]. We extend this result to the full modal language in a later section.
Lemma 4.13.
For any 1-isolator $\varphi$, if $\varphi$ is $\mathbf{S5[Con,Ground]}$-consistent, then there is some $\star$ and some $w$ such that $\mathbf{FIX},w\Vdash_{V^{\star}}\varphi$.
Proof.
Suppose $\varphi$ is a 1-isolator, i.e., $\varphi=\varepsilon(x)\wedge\iota(x)$ where $\varepsilon(x)$ is an extensional 1-isolator and $\iota(x)$ is an intensional 1-isolator. $\mathbf{S5[Con,Ground]}\vdash\varepsilon(x)\to\lozenge\varepsilon(x)$. Hence $\iota(x)$ has $\lozenge\varepsilon(x)$ as a conjunct.
We now break into cases depending on which extensional 1-isolator $\varepsilon(x)$ is. In each case we will break into subcases, checking each intensional 1-isolator that has $\lozenge\varepsilon(x)$ as a conjunct. In each subcase we find an assignment $\star$ that satisfies $\varphi$.
Case 1: Suppose that $\varepsilon(x)$ is $T(x)$. We must check the following intensional 1-isolators:
  1. (i)
    $\iota(x)=\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$. Let $x^{\star}=\tau$. Then there is some $w$ such that $\mathbf{FIX},w\Vdash_{V^{\star}}\varepsilon(x)\wedge\iota(x)$.
  2. (ii)
    $\iota(x):=\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$. Let $x^{\star}=\gamma$. Then there is some $w$ such that $\mathbf{FIX},w\Vdash_{V^{\star}}\varepsilon(x)\wedge\iota(x)$.
  3. (iii)
    $\iota(x):=\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\neg\lozenge N(x)$. Let $x^{\star}=0=0$. Then for any $w$, $\mathbf{FIX},w\Vdash_{V^{\star}}\varepsilon(x)\wedge\iota(x)$.
We leave Cases 2 ($\varepsilon(x)$ is $F(x)$) and 3 ($\varepsilon(x)$ is $N(x)$) to the reader. ∎
Theorem 4.14.
For all 1-formulas $\varphi\in\mathcal{L}_{\Box}$, the following are equivalent:
  1. (i)
    $\mathbf{FIX}\Vdash\varphi$
  2. (ii)
    $\mathbf{S5[Con,Ground]}\vdash\varphi$
Proof.
The soundness direction is Proposition 3.8. Let’s focus on completeness. Suppose that $\mathbf{S5[Con,Ground]}\nvdash\varphi$. Then, by classical logic, $\mathbf{S5[Con,Ground]}+\neg\varphi$ is consistent. By Theorem 4.9,
$\mathbf{S5[Con,Ground]}\vdash\neg\varphi\leftrightarrow(\psi_{1}\vee\dots\vee\psi_{k})$
where each $\psi_{i}$ is a 1-isolator. By classical logic, at least one of these 1-isolators $\psi_{i}$ is consistent. By Lemma 4.13, we infer that there is some $\star$ such that $\mathbf{FIX}\Vdash_{V^{\star}}\psi_{i}$. Hence $\mathbf{FIX}\Vdash_{V^{\star}}\neg\varphi$, so $\mathbf{FIX}\nVdash\varphi$. ∎
Proposition 4.15.
Every $\mathbf{FIX}$-definable $A\subseteq\mathcal{L}_{T}$ is definable by a (possibly empty) disjunction of intensional 1-isolators for $\mathbf{S5[Con,Ground]}$.
Proof.
Every $\mathbf{FIX}$-definable $A\subseteq\mathcal{L}_{T}$ is definable by an intensional 1-formula. Indeed, if $\varphi$ defines $A$, then this means that exactly the elements of $A$ satisfy $\varphi$ at every world. So $\Box\varphi$ also defines $A$. The result follows since every intensional 1-formula is equivalent to a disjunction of intensional 1-isolators (see Theorem 4.9). ∎
Corollary 4.16.
The number of $\mathbf{FIX}$-definable sets is exactly $64$.
Proof.
There are exactly $6$ intensional 1-isolators for $\mathbf{S5[Con,Ground]}$. $642^{6}=$ 64. The result then follows from Proposition 4.15. ∎
4.4. Descriptive complexity
In this subsection we will finally fulfill our promise to prove the following result:
Theorem 4.17.
The set of intrinsic sentences is not $\mathbf{FIX}$-definable.
There are multiple ways of verifying this, and we will present two. One could check all 64 single variable definitions, although this would be tedious. Luckily, there are more conceptual arguments.
Every $\mathbf{FIX}$-definable is definable by an intensional 1-formula. This is because if $\varphi$ $\mathbf{FIX}$-defines a set $\mathcal{A}$, then so does $\Box\varphi$. We have already seen that each intensional 1-formula is $\mathbf{S5[Con,Ground]}$-equivalent to a disjunction of intensional 1-isolators. So, pushing quantifiers in, we see that every modal definition is equivalent to a formula of the following form:
($\big{(}{}_{\lozenge}^{\Box}{}_{\neg}^{+}{}_{F}^{T}(x)\wedge\dots\wedge{}_{\lozenge}^{\Box}{}_{\neg}^{+}{}_{F}^{T}(x)\big{)}\vee\dots\vee({}_{\lozenge}^{\Box}{}_{\neg}^{+}{}_{F}^{T}(x)\wedge\dots\wedge{}_{\lozenge}^{\Box}{}_{\neg}^{+}{}_{F}^{T}(x)\big{)}$
Let’s say that any such formula is in basic form.
Definition 4.18.
Let $\mathsf{Fix}$ be an $\mathcal{L}_{2}$ formula defining the set of fixed points. We define a translation that maps $\star:\mathcal{L}_{\Box}$ formulas in basic form to $\mathcal{L}_{2}$ formulas.
$(A\vee\dots\vee A)^{\star}:=A^{\star}\vee\dots\vee A^{\star}$
$(\Box_{F}^{T}(x))^{\star}:=\forall X\big{(}\mathsf{Fix}(X)\to\ulcorner_{F}^{T}(x)^{\star}\urcorner\in X\big{)}$
($(\Box\neg_{F}^{T}(x))^{\star}:=\forall X\big{(}\mathsf{Fix}(X)\to\ulcorner_{F}^{T}(x)^{\star}\urcorner\notin X\big{)}$
$(\lozenge_{F}^{T}(x))^{\star}:=\exists X\big{(}\mathsf{Fix}(X)\wedge\ulcorner_{F}^{T}(x)^{\star}\urcorner\in X\big{)}$
($(\lozenge\neg_{F}^{T}(x))^{\star}:=\exists X\big{(}\mathsf{Fix}(X)\wedge\ulcorner_{F}^{T}(x)^{\star}\urcorner\notin X\big{)}$
$T(x)^{\star}=x$
$F(x)^{\star}=Neg(x^{\star})$
Lemma 4.19.
If $\varphi$ modally defines $A$, then $\varphi^{\star}$ defines $A$ over the standard interpretation of second-order arithmetic.
Proof.
This follows directly by analyzing the semantics of the modal language. ∎
Theorem 4.20.
Every modally definable set is $\mathsf{Boolean}\text{-}\Sigma^{1}_{1}$-definable in the language of second-order arithmetic.
Proof.
The set of fixed points is $\Delta^{1}_{1}$-definable (see [burgess1986truth, §6]). In particular, we can choose a $\Sigma^{1}_{1}$ formula $\mathsf{Fix}$ defining the set of fixed points. ∎
The desired indefinability result then follows from the following theorem:
Theorem 4.21 ([burgess1986truth]).
The set of intrinsic sentences is $\Sigma^{1}_{1}$-complete in a $\Pi^{1}_{1}$ parameter.
Corollary 4.22.
The set of intrinsic sentences is not modally definable.
Proof.
No $\Sigma^{1}_{1}$-complete in a $\Pi^{1}_{1}$ parameter set admits a $\mathsf{Boolean}\text{-}\Sigma^{1}_{1}$ definition. ∎
4.5. Indiscernibility
The previous result demonstrates the non-definability of a great number of sets, but it is a coarse tool. For instance, descriptive complexity considerations do not yield the non-definability of the arithmetic truths, since they form a $\Delta^{1}_{1}$ set, but the the arithmetic truths are not definable.
We will now prove the indefinability of the intrinsic sentences in another way. Informally, sets $\mathcal{A}$ and $\mathcal{B}$ are $\mathbf{FIX}$-indiscernible just in case every formula that applies to the $\mathcal{A}$s applies to the $\mathcal{B}$s and vice-versa. We will prove that the intrinsic sentences and the inevitable sentences are $\mathbf{FIX}$-indiscernible, whence the set of intrinsic sentences—the smaller of the two sets—cannot be definable.
Definition 4.23.
$\mathcal{A}$ and $\mathcal{B}$ are $\mathbf{FIX}$-indiscernible if for every $\varphi(x)\in\mathcal{L}_{\Box}$, TFAE:
  1. (1)
    For all $\star$ with $x^{\star}\in\mathcal{A}$, $\mathbf{FIX}\Vdash_{\mathbf{V^{\star}}}\varphi(x)$.
  2. (2)
    For all $\star$ with $x^{\star}\in\mathcal{B}$, $\mathbf{FIX}\Vdash_{\mathbf{V^{\star}}}\varphi(x)$.
Theorem 4.24.
The set of intrinsic sentences and the set of inevitable sentences are $\mathbf{FIX}$-indiscernible.
Proof.
Let $\mathcal{A}$ be the set of intrinsic and $\mathcal{B}$ the set of inevitable sentences. Then (2) implies (1) (Definition 4.23) is trivial.
It suffices to check that (1) implies (2). So assume (1), i.e.,
$\oplus\quad\text{For all $\star$ with $x^{\star}$ intrinsic, $\mathbf{FIX}\Vdash_{\mathbf{V^{\star}}}\varphi(x)$}.$
We may assume that $\varphi$ is an intensional formula (if it is not, just stick a $\Box$ in front of it). We may then assume that $\varphi$ is in normal form, i.e., $\varphi$ is a disjunction of intensional 1-isolators.
Now the intensional 1-isolators are the following six formulas:
  1. (i)
    $\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
  2. (ii)
    $\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
  3. (iii)
    $\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
  4. (iv)
    $\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\neg\lozenge N(x)$
  5. (v)
    $\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\neg\lozenge N(x)$
  6. (vi)
    $\neg\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
Note that the set of inevitable sentences is defined by the disjunction $(ii)\vee(iii)\vee(iv)\vee(v)$.
Suppose, for the sake of contradiction, that (2) fails. Then there is some $\star$ such that $y^{\star}=\psi$ is inevitable and that $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(y)$. So $\varphi$ must be missing one of the disjuncts required to define the inevitable sentences, i.e., (ii)–(iv). Let’s break into cases.
Case 1: $\varphi$ is missing disjunct (ii). But $\gamma$ is intrinsic and $ii(\gamma)$. So if $x^{\star}=\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
Case 2: $\varphi$ is missing disjunct (iii). But $\neg\gamma$ is intrinsic and $iii(\neg\gamma)$. So if $x^{\star}=\neg\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
Case 3: $\varphi$ is missing disjunct (iv). But $0=0$ is intrinsic and $iv(0=0)$. So if $x^{\star}=0=0$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
Case 4: $\varphi$ is missing disjunct (v). But $0=1$ is intrinsic and $v(0=1)$. So if $x^{\star}=0=1$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$. ∎
Remark 4.25.
If $\mathcal{A}$ and $\mathcal{B}$ are $\mathbf{FIX}$-indiscernible and $\mathcal{A}\subset\mathcal{B}$, then $\mathcal{A}$ is not $\mathbf{FIX}$-definable. Hence Theorem 4.24 entails that the set of intrinsic sentences is not $\mathbf{FIX}$-definable.
5. Multivariable Definability
The logic and combinatorics become more involved—but also more interesting—when there are many variables. 1-formulas are used to define sets of sentences. 2-formulas (i.e., formulas with two distinct variables) are used to define binary relations on sentences. Likewise, $n$-formulas (i.e., formulas with $n$ distinct variables) are used to define $n$-ary relations on sentences.
Definition 5.1.
An $\mathcal{L}_{\Box}$ formula $\varphi(x_{1},\dots,x_{n})$ defines $\mathcal{A}$ over $\mathbf{FIX}$ just in case for every $\mathcal{L}_{T}$-realization $\star$:
$\mathbf{FIX}\Vdash_{V^{\star}}\varphi(x_{1},\dots,x_{n})\Longleftrightarrow\langle x_{1}^{\star},\dots,x_{n}^{\star}\rangle\in\mathcal{A}.$
5.1. Counting Formulas
Let’s begin with some relatively easy counting problems. We will turn to more difficult counting problems in a later subsection.
5.1.1. The Bare Modal System
Carnap’s normal form theorems generalize to $n$ variables. We summarize the important information for $\mathbf{S5}$ with the following bullets:
  • The extensional $n$-isolators are the formulas of the form:
    $\mathrel{\begin{subarray}{c}B\\ T_{only}\\ F_{only}\\ N\end{subarray}}(x_{1})\wedge\dots\wedge\mathrel{\begin{subarray}{c}B\\ T_{only}\\ F_{only}\\ N\end{subarray}}(x_{n})$
  • The intensional pre-n-isolators are the formulas of the form:
    ¬${}^{+}_{\neg}\lozenge\varphi_{1}\wedge\dots\wedge^{+}_{\neg}\lozenge\varphi_{4^{n}}$
    where $\varphi_{1},\dots,\varphi_{4^{n}}$ are the extensional $n$-isolators.
  • $\mathbf{S5}$ refutes exactly one of these, namely:
    $\neg\lozenge\varphi_{1}\wedge\dots\wedge\neg\lozenge\varphi_{4^{n}}$
    and the rest are the intensional $n$-isolators.
  • The $n$-isolators are the consistent conjunctions $\varphi\wedge\psi$ for $\varphi$ an extensional $n$-isolator and $\psi$ an intensional $n$-isolator.
With this terminology on board, we can state the general version of Carnap’s theorem:
Theorem 5.2 (essentially [carnap1946modalities]).
For $T$ extending $\mathbf{S5}$:
  1. (1)
    Every extensional $n$-formula is $T$ equivalent to a disjunction of extensional $n$-isolators for $T$.
  2. (2)
    Every intensional $n$-formula is $T$ equivalent to a disjunction of intensional $n$-isolators for $T$.
  3. (3)
    Every $n$-formula is $T$ equivalent to a disjunction of $n$-isolators for $T$.
Let’s solve a counting problem.
Proposition 5.3.
Up to $\mathbf{S5}$-provable equivalence, the number of $n$-formulas is exactly: $2^{(4^{n}2^{(4^{n}-1)})}$.
Proof.
$\mathbf{S5}$ has $4^{n}$ extensional $n$-isolators and $12^{(4^{n})}-$ 1 intensional $n$-isolators.
Each extensional $n$-isolator is inconsistent with the intensional 1-isolators claiming that it is impossible, and consistent otherwise.
So fixing a particular extensional $n$-isolator $\varepsilon$, there are $2^{(4^{n}-1)}$ many intensional $n$-isolators with $\lozenge\varepsilon$ as a conjunct.
So there are $4^{n}\times 2^{(4^{n}-1)}$ many $n$-isolators. So the number of formulas up to provable equivalence is $2^{(4^{n}2^{(4^{n}-1)})}$. ∎
5.1.2. With the Consistency Axiom
Let’s turn out attention to $\mathbf{S5[Con]}$. We summarize the important information in the following bullets:
  • The extensional $n$-isolators are the formulas of the form:
    $\mathrel{\begin{subarray}{c}T\\ F\\ N\end{subarray}}(x_{1})\wedge\dots\wedge\mathrel{\begin{subarray}{c}T\\ F\\ N\end{subarray}}(x_{n})$
  • The intensional pre-n-isolators are the formulas of the form:
    ¬${}^{+}_{\neg}\lozenge\varphi_{1}\wedge\dots\wedge^{+}_{\neg}\lozenge\varphi_{3^{n}}$
    where $\varphi_{1},\dots,\varphi_{3^{n}}$ are the extensional $n$-isolators.
  • $\mathbf{S5[Con]}$ refutes exactly one of these, namely:
    $\neg\lozenge\varphi_{1}\wedge\dots\neg\lozenge\varphi_{3^{n}}$
    and the rest are the intensional $n$-isolators.
  • The $n$-isolators are the consistent conjunctions $\varphi\wedge\psi$ for $\varphi$ an extensional $n$-isolator and $\psi$ an intensional $n$-isolator.
Proposition 5.4.
Up to $\mathbf{S5[Con]}$-provable equivalence, the number of $n$-formulas is exactly: $2^{(3^{n}2^{(3^{n}-1)})}$.
Proof.
$\mathbf{S5[Con]}$ has $3^{n}$ extensional $n$-isolators and $12^{(3^{n})}-$ 1 intensional $n$-isolators.
Each extensional $n$-isolator is inconsistent with the intensional 1-isolators claiming that it is impossible, and consistent otherwise.
So fixing a particular extensional $n$-isolator $\varepsilon$, there are $2^{(3^{n}-1)}$ many intensional $n$-isolators with $\lozenge\varepsilon$ as a conjunct.
So there are $3^{n}\times 2^{(3^{n}-1)}$ many $n$-isolators. So the number of formulas up to provable equivalence is $2^{(3^{n}2^{(3^{n}-1)})}$. ∎
5.2. Matrices
The combinatorics of definability is much harder when $n>1$. In this section, for the sake of developing intuitions, we focus on the 2-variable case. Let’s begin by considering this matrix:
TT TF TN
FT FF FN
NT NF NN
Table 1. All extensional 2-isolators.
We identify each cell of the matrix with an extensional 2-isolator. In particular, we identify the cell AB with the formula $A(x_{1})\wedge B(x_{2})$. For instance, we identify the cell TN with the formula $T(x_{1})\wedge N(x_{2})$. This yields an association of extensional 2-isolators with pairs of numbers, via Table 2. We will often consider subsets $\mathcal{S}$ of the $3\times 3$ matrix and we move freely between writing as though $\mathcal{S}$ contains NN, $N(x_{1})\wedge N(x_{2})$, and (3,3), trusting that no confusion will arise.
1,1 1,2 1,3
2,1 2,2 2,3
3,1 3,2 3,3
Table 2. Standard numbering of the $3\times 3$ matrix.
Each intensional 2-isolator for $\mathbf{S5[Con]}$ corresponds to a subset of the $3\times 3$ matrix. In particular, each intensional 2-isolator has the form:
$\bigwedge\{\lozenge\varphi\mid\varphi\in\mathcal{S}\}\wedge\bigwedge\{\neg\lozenge\varphi\mid\varphi\notin\mathcal{S}\}$
where $\mathcal{S}$ is some subset of the matrix.
However, it is not the case that each subset $\mathcal{S}$ of the matrix yields an intensional 2-isolator for $\mathbf{S5[Con]}$. The issue is that some of the subsets of the matrix correspond to $\mathbf{S5[Con]}$-refutable formulas, and intensional 2-isolators are consistent by definition. For instance, letting $\mathcal{S}$ be the empty set yields the formula:
$\bigwedge\{\neg\lozenge\varphi\mid\varphi\notin\emptyset\}$
which is refuted by $\mathbf{S5[Con]}$, whence it is not an intensional 2-isolator.
The crucial question for us is to identify which subsets of the matrix actually reflects the intensional properties of $\mathcal{L}_{T}$ sentences. Let’s review some examples.
Consider Table 3, Table 4, Table 5, and Table 6. Each of these images exhibits a subset of the $3\times 3$ matrix. In each case, consider the following question: Is there a variable assignment $\star$ such that $x_{1}^{\star}$ and $x_{2}^{\star}$ can attain all and only these combinations of values? Put another way, given a subset $\mathcal{S}$ of the $3\times 3$ matrix, is there a variable assignment $\star$ such that $\mathbf{FIX}_{V^{\star}}\Vdash\bigwedge\{\lozenge\varphi\mid\varphi\in\mathcal{S}\}\wedge\bigwedge\{\neg\lozenge\varphi\mid\varphi\notin\mathcal{S}\}$?
TT TF TN
Table 3. A proper subset of the extensional 2-isolators.
TT TF
Table 4. A proper subset of the extensional 2-isolators.
TT
FF
NN
Table 5. A proper subset of the extensional 2-isolators.
TT TN
FN
NT
Table 6. A proper subset of the extensional 2-isolators.
The answers are given here:
Table 3: Yes. Let $x_{1}^{\star}$ be a grounded truth (e.g., $0=0$) and $x_{2}^{\star}$ be a truth-teller.
Table 4: No. Note that the first two columns are non-empty but the third is empty. Thus, the axiom $\mathbf{Ground}$
$\big{(}\lozenge T(x)\wedge\lozenge F(x)\big{)}\to\lozenge N(x)$
rules out this possibility, and $\mathbf{FIX}\Vdash\mathbf{Ground}$.
Table 5: Yes. Let $x_{1}^{\star}$ and $x_{2}^{\star}$ be the same truth-teller.
Table 6: No. However, one cannot see that there is no such assignment merely deploying the principles codified in $\mathbf{S5[Con,Ground]}$. Indeed, we must introduce the following axiom:
$\mathbf{Min_{2}}:=\big{(}\lozenge N(x)\wedge\lozenge N(y)\big{)}\to\lozenge\big{(}N(x)\wedge N(y)\big{)}$
Even though $\mathbf{S5[Con,Ground]}\nvdash\mathbf{Min_{2}}$, it is still the case that $\mathbf{FIX}\Vdash\mathbf{Min_{2}}$ (by Remark 2.4). Note that the axiom $\mathbf{Min_{2}}$ rules out Table 6 since the third row is non-empty and the third column is non-empty but the corner element NN is missing.
This last example identifies the only constraint on intensional 2-isolators that is not codified in the $\mathbf{S5[Con,Ground]}$ axioms.
5.3. The Prime Conditions
For the full completeness theorem we need new axioms:
$\mathbf{Min_{n}}:=\big{(}\lozenge N(x_{1})\wedge\dots\wedge\lozenge N(x_{n})\big{)}\to\lozenge\big{(}N(x_{1})\wedge\dots\wedge N(x_{n})\big{)}$
Remark 5.5.
For $n>k$, $\mathbf{Min_{n}}\vdash\mathbf{Min_{k}}$. To see this, just pick an instance of $\mathbf{Min_{n}}$ wherein some variables are repeated. For instance, to see that $\mathbf{Min_{3}}\vdash\mathbf{Min_{2}}$ note that the following instance of $\mathbf{Min_{3}}$:
$\big{(}\lozenge N(x_{1})\wedge N(x_{1})\wedge\lozenge N(x_{2})\big{)}\to\lozenge\big{(}N(x_{1})\wedge N(x_{1})\wedge N(x_{2})\big{)}$
is (logically equivalent to) an instance of $\mathbf{Min_{2}}$.
In the previous subsection we identified intensional $2$-isolators with subsets of the $3\times 3$ matrix. For the $n$-variable case, we must turn from matrices to tensors, which are generalizations of matrices beyond two dimensions. In particular, will be interested in $3^{n}$ tensors, i.e., $\underbrace{3\times\ldots\times 3}\limits_{\mbox{\scriptsize$n$ times}}$ matrices.
Let $[3]^{n}$ be the set of triples from the set $\{1,2,3\}$. This is the set of positions in a $3^{n}$ tensor. We can no longer speak only of rows and columns. Instead, we speak in terms of slices. The rows are the 1-slices, the columns are the 2-slices, and so on. How do we speak of the first row, the second row, the third column, etc.? The 1st row is the 1st layer in the 1st slice, the third column is the 3rd layer in the second slice, and so on. We introduce the following notation.
Definition 5.6.
For $k\leq 3$ and $j\leq n$, the $k^{th}$-layer of the $j^{th}$ slice (written $j_{k}$) is
$\{a\in[3^{n}]\mid\exists b_{1},\dots,\exists b_{n-1}\;a=\langle b_{1},\dots b_{j-1},k,b_{j},\dots,b_{n-1}\rangle\}$
Note that as $j_{k}$ is defined, it is a set of positions in the $3^{n}$ tensor. That is, so far we have still only discussed positions. The tensors we are interested in are 0-1 valued tensors, i.e., assignments of 0 and 1 to these positions. It is easiest to think of these as just subsets of $[3]^{n}$, and that is how we will proceed.
We now introduce a definition that will be crucial to the proofs of the remaining theorems:
Definition 5.7.
A tensor $\mathcal{S}\subseteq[3]^{n}$ meets the prime conditions if:
  1. (1)
    $\mathcal{S}$ is non-empty.
  2. (2)
    For all $j\leq n$, if $\mathcal{S}$ intersects $j_{1}$ and $j_{2}$, then $\mathcal{S}$ intersects $j_{3}$.
  3. (3)
    For all $l\leq n$, for any $t$ slices $s^{1},\dots,s^{t}$, if $\mathcal{S}$ intersects $s^{i}_{3}$ for each $i\leq t$, then $\mathcal{S}$ intersects $s^{1}_{3}\cap\dots\cap s^{t}_{3}$.
Remark 5.8.
In the case $n=2$, the prime conditions can be stated very simply (note that when we refer to row 1, column 2, etc. we have in mind Table 1). We state them here to aid the reader’s intuition. A set $\mathcal{S}$ of extensional 2-isolators meets the prime conditions if:
  1. (1)
    $\mathcal{S}$ is non-empty.
  2. (2)
    1. (a)
      If $\mathcal{S}$ intersects row 1 and row 2, it intersects row 3.
    2. (b)
      If $\mathcal{S}$ intersects column 1 and column 2, it intersects column 3.
  3. (3)
    If $\mathcal{S}$ intersects row 3 and column 3, it contains (3,3).
5.4. The Main Lemma
The main technical lemma of this paper connects the prime conditions with $\mathbf{FIX}$-satisfiability. That is, this lemma shows that those sets that satisfy the prime conditions are exactly the sets that reflect the intensional properties of $\mathcal{L}_{T}$ sentences.
Lemma 5.9.
For $\mathcal{S}\subseteq[3]^{n}$, TFAE:
  1. (1)
    $\mathcal{S}$ meets the prime conditions.
  2. (2)
    There is an $\mathcal{L}_{T}$ realization $\star$ such that
    $\mathbf{FIX}\Vdash_{V^{\star}}\bigwedge\{\lozenge\varphi\mid\varphi\in\mathcal{S}\}\wedge\bigwedge\{\neg\lozenge\varphi\mid\varphi\notin\mathcal{S}\}.$
Proof.
(2) implies (1): This follows from soundness, i.e.,
$\mathbf{FIX}\Vdash\mathbf{S5[Con,Ground,Min_{n}]}.$
See Remark 2.1 and Remark 2.4.
(1) implies (2): Suppose that $\mathcal{S}$ meets the prime conditions. There are two cases to consider.
  1. (I)
    $\mathcal{S}$ does not contain $(3,\dots,3)$.
  2. (II)
    $\mathcal{S}$ does contain $(3,\dots,3)$.
Case (I): There are very few ways a set can meet the prime conditions and omit (3,…,3). The important thing to note is that:
Claim 1.
For some $j\leq n$ and some $i\leq 2$, $\mathcal{S}$ is contained in the $i$th layer of the $j$th slice.
Proof.
To see that the claim holds, suppose that for all $j\leq n$, for all $i\leq 2$, $\mathcal{S}$ is not contained in the $i$th layer of the $j$th slice. Since $\mathcal{S}$ meets the prime conditions, $\mathcal{S}$ is non-empty. Thus, for every $j$, $\mathcal{S}$ must intersect some layer of the $j$th slice. We reason as follows:
  • $\mathcal{S}$ is not contained in the first layer of the $j$th slice, so $\mathcal{S}$ intersects either the second or third layer of the $j$th slice.
  • $\mathcal{S}$ is not contained in the second layer of the $j$th slice, so $\mathcal{S}$ intersects either the first or third layer of the $j$th slice.
It follows that either $\mathcal{S}$ intersects $j_{3}$ or it intersects both $j_{1}$ and $j_{2}$. In the latter case, the prime conditions imply that $\mathcal{S}$ intersects $j_{3}$. So either way, $\mathcal{S}$ intersects $j_{3}$.
So, for every $j$, $\mathcal{S}$ intersects $j_{3}$. The prime conditions imply that $\mathcal{S}$ contains $(3,\dots,3)$, a contradiction. This completes the proof of the claim. ∎
With the claim on board we now set about finishing Case (I) by induction. The base case is $n=2$.
Base Case: We can split into the subcases provided by the claim ($\mathcal{S}$ is contained either in the first row, the second row, the first column, or the second column). Suppose for simplicity’s sake that $\mathcal{S}$ is contained in the first row. There are only so many sets contained in the first row:
  1. (i)
    $\mathcal{S}$ is the first row. Let $A^{\star}$ be $0=0$ and $B^{\star}$ be $\tau$.
  2. (ii)
    $\mathcal{S}$ contains just the leftmost and the rightmost elements of the first row. Let $A^{\star}$ be $0=0$ and $B^{\star}$ be $\gamma$.
  3. (iii)
    $\mathcal{S}$ contains the two rightmost elements of the first row. Let $A^{\star}$ be $0=0$ and $B^{\star}$ be $\neg\gamma$.
  4. (iv)
    $\mathcal{S}$ is just the leftmost element. Let $A^{\star}$ be 0=0 and $B^{\star}$ be 0=0.
  5. (v)
    $\mathcal{S}$ is just the middle element. Let $A^{\star}$ be 0=0 and $B^{\star}$ be 0=1.
  6. (vi)
    $\mathcal{S}$ is just the rightmost element. Let $A^{\star}$ be 0=0 and $B^{\star}$ be $\lambda$.
By the prime conditions, $\mathcal{S}$ cannot contain just the two leftmost elements of the first row, so these cases are exhaustive.
The reasoning is similar if $\mathcal{S}$ is contained in the second row, first column, or second column. We leave these cases to the reader.
Induction step: Assume the induction hypothesis: For any $\mathcal{S}\subseteq[3]^{n}$ meeting the prime conditions, there is an appropriate realization $\star$. Now let $\mathcal{T}\subseteq[3]^{n+1}$ meet the prime conditions. By Claim 1, $\mathcal{T}$ is contained in the $i$th layer of the $j$th slice. The key observation is that the $i$th layer of the $j$th slice is a $3^{n}$ tensor. More precisely, the $j$th conjunct is the same in all the entries in $j_{i}$, so ignoring this conjunct in all entries yields a $3^{n}$ tensor $\mathcal{T}\restriction[3]^{n}$. Apply the induction hypothesis to get an $\mathcal{L}_{T}$ realization $\star$ such that
$\mathbf{FIX}\Vdash_{V^{\star}}\bigwedge\{\lozenge\varphi\mid\varphi\in\mathcal{T}\restriction[3]^{n}\}\wedge\bigwedge\{\neg\lozenge\varphi\mid\varphi\notin\mathcal{T}\restriction[3]^{n}\}.$
To extend $\star$ to a realization $\sharp$ for $\mathcal{T}$, just let
$x_{j}^{\sharp}=\begin{cases}0=0&\text{ if }i=1\\ 0=1&\text{ if }i=2\\ \lambda&\text{ if }i=3\end{cases}$
Case (II): Assume that $(3,\dots,3)\in\mathcal{S}$. We must define sentences $A^{1},\dots,A^{n}$. Pick some $k$ such that $2^{k}>3^{n}$. Take a set $\{\tau_{1},\dots,\tau_{k}\}$ of independent truth-tellers (see Remark 2.2). There are $2^{k}$-many formulas of the form:
${}^{+}_{\neg}\tau_{1}\wedge\dots\wedge^{+}_{\neg}\tau_{k}$
Fix some enumeration $C_{1},\dots,C_{2^{k}}$ of these formulas. We then define the formulas $A^{1}_{1},\dots,A^{1}_{3^{n}},\dots,A^{n}_{1},\dots,A^{n}_{3^{n}}$ so that $A^{i}_{j}$ is a conditional of the form $C_{j}\to\theta$ where $\theta\in\{0=0,0=1,\lambda\}$.
More precisely, we proceed as follows. We fix a numbering $\Phi$ of the cells in the $3^{n}$ tensor. Then $\Phi(j)$ is some formula $\varphi_{1}(x_{1})\wedge\dots\wedge\varphi_{n}(x_{n})$. Let $\Phi(j)_{i}$ be the $i$th conjunct of $\Phi(j)$. We then define:
$A^{i}_{j}=C_{j}\to\begin{cases}0=0&\text{ if }\Phi(j)_{i}=T(x_{i})\\ 0=1&\text{ if }\Phi(j)_{i}=F(x_{i})\\ \lambda&\text{ if }\Phi(j)_{i}=N(x_{i})\end{cases}$
We define $\mathcal{S}^{\Phi}:=\{k\leq 3^{n}\mid\Phi(k)\in\mathcal{S}\}$.
We define $A^{i}$ as follows:
$A^{i}\equiv\bigwedge\{A^{i}_{j}\mid j\in\mathcal{S}^{\Phi}\}\wedge\bigwedge\{C_{j}\to\lambda\mid j\in\{1,\dots,2^{k}\}-\mathcal{S}^{\Phi}\}$
Let $\star$ be a variable assignment such that $x_{i}^{\star}=A^{i}$ for each $i\leq n$.
Claim 2.
If $\varphi\in\mathcal{S}$, then $\mathbf{FIX}\Vdash_{V^{\star}}\lozenge\varphi$.
Proof.
Suppose that $\varphi\in\mathcal{S}$. Note that $\varphi$ is a conjunction of the form $\varphi_{1}(x_{1})\wedge\dots\wedge\varphi_{n}(x_{n})$. Note also that $\Phi^{-1}(\varphi)$ is some number $k\leq 3^{n}$. Since $C_{k}$ is a Boolean statement about independent truth-tellers, there is some $w$ such that $\mathbb{N},w\vDash_{\mathsf{K3}}C_{k}$ (see Remark 2.2).
Let $i\leq n$. Note that $A^{i}_{k}$ is a conditional of the form $C_{k}\to\theta$.
$A^{i}$ is a conjunction of conditionals with antecedent $C_{j}$ such that for $j\neq k$, $\mathbb{N},w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}C_{j}$. Which is to say that for all conjuncts $A^{i}_{j}$ of $A^{i}$ except for $A^{i}_{k}$, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{j}$. Thus, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{k}$. Since $\mathbb{N},w\vDash_{\mathsf{K3}}C_{k}$, we infer that, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta$.
We now want to prove that $\mathbf{FIX},w\Vdash_{V^{\star}}\varphi_{i}$, where $\varphi_{i}$ is the $i$th conjunct of $\varphi$. Let’s break into cases depending on which formula $\varphi_{i}$ is.
Case 1: Suppose $\varphi_{i}$ is $T(x_{i})$, i.e, $\varphi$ is in the 1st layer of the $i$th slice. Since $k=\Phi^{-1}(\varphi)$, $\Phi(j)_{i}=T(x_{i})$. So, $\theta$ is $0=0$. So $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}$. So $\mathbb{N},w\vDash_{\mathsf{K3}}\mathsf{True}(\ulcorner A^{i}\urcorner)$. Since $x_{i}^{\star}=A^{i}$, $\mathbf{FIX},w\Vdash_{V^{\star}}T(x_{i})$.
Case 2: Suppose $\varphi_{i}$ is $F(x_{i})$, i.e, $\varphi$ is in the 2nd layer of the $i$th slice. Since $k=\Phi^{-1}(\varphi)$, $\Phi(j)_{i}=F(x_{i})$. So, $\theta$ is $0=1$. So $\mathbb{N},w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}A^{i}$. So $\mathbb{N},w\vDash_{\mathsf{K3}}\mathsf{True}(\ulcorner\neg A^{i}\urcorner)$. Since $x_{i}^{\star}=A^{i}$, $\mathbf{FIX},w\Vdash_{V^{\star}}F(x_{i})$.
Case 3: Suppose $\varphi_{i}$ is $N(x_{i})$, i.e, $\varphi$ is in the 3rd layer of the $i$th slice. Since $k=\Phi^{-1}(\varphi)$, $\Phi(j)_{i}=N(x_{i})$. So, $\theta$ is $\lambda$. So $\mathbb{N},w\uparrow_{\mathsf{K3}}A^{i}$. So $\mathbb{N},w\nVdash_{\mathsf{K3}}\mathsf{True}(\ulcorner A^{i}\urcorner)$ and $\mathbb{N},w\nVdash_{\mathsf{K3}}\mathsf{True}(\ulcorner\neg A^{i}\urcorner)$. Since $x_{i}^{\star}=A^{i}$, $\mathbf{FIX},w\Vdash_{V^{\star}}N(x_{i})$.
So $\mathbf{FIX},w\Vdash_{V^{\star}}\varphi$. So for all $v\in\mathbf{FIX}$, $\mathbf{FIX},v\Vdash_{V^{\star}}\lozenge\varphi$. That is, $\mathbf{FIX}\Vdash_{V^{\star}}\lozenge\varphi$. ∎
Claim 3.
If $\varphi\notin\mathcal{S}$, then $\mathbf{FIX}\nVdash_{V^{\star}}\lozenge\varphi$.
Proof.
We prove the contrapositive. Suppose $\mathbf{FIX}\Vdash_{V^{\star}}\lozenge\varphi$. Then there is some $w$ such that $\mathbf{FIX},w\Vdash_{V^{\star}}\varphi$. We now break into cases depending on whether any of the $C_{s}$ is true at $w$.
Case 1: Suppose $\mathbb{N},w\uparrow_{\mathsf{K3}}C_{s}$ for all $s\leq 16$. We claim that $A^{1},\dots,A^{n}$ all have value neither in $w$.
To see this, note that $A^{i}$ is a conjunction of conditionals each of which has an antecedent $C_{s}$, which has value neither. A conditional with a neither antecedent cannot be false, so $A^{i}$ cannot be false. $A_{i}$ is true only if all its conjuncts are true. Since $2^{k}>3^{n}\geq|\mathcal{S}^{\Phi}|$, there is some $j\leq 2^{k}$ such that $j\notin\mathcal{S}^{\Phi}$. But then by definition of $A^{i}$, $C_{j}\to\lambda$ is a conjunct of $A^{i}$, and this has value neither.
So $\varphi$ is the formula $N(x_{1})\wedge\dots\wedge N(x_{n})$. Since we are working in Case (II), we have assumed that this formula belongs to $\mathcal{S}$.
Case 2: Suppose $\mathbb{N},w\vDash_{\mathsf{K3}}C_{s}$ for exactly one $s$. We split into subcases depending on whether $s>3^{n}$ or $s\leq 3^{n}$.
Subcase 2.1: Suppose $s>3^{n}$. $A^{i}$ is a conjunction of conditionals of the form $C_{j}\to\theta^{A}_{i}$ such that for $j\neq s$, $\mathbb{N},w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}C_{j}$, whence $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{j}$. Accordingly, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta$, where $\theta$ is the consequent of $A^{i}_{s}$. Since $s>3^{n}$, $\theta$ is $\lambda$. Hence, $\mathbb{N},w\uparrow_{\mathsf{K3}}A^{i}$. Hence, $\mathbb{N},w\nVdash_{\mathsf{K3}}\mathsf{True}(\ulcorner A^{i}\urcorner)$ and $\mathbb{N},w\nVdash_{\mathsf{K3}}\mathsf{True}(\ulcorner\neg A^{i}\urcorner)$. Since $x_{i}^{\star}=A^{i}$, $\mathbf{FIX},w\Vdash_{V^{\star}}N(x_{i})$.
Thus, $\varphi$ is $N(x_{1})\wedge\dots\wedge N(x_{n})$, i.e., (3,…, 3). Once again, since we are working in Case (II), we have assumed that this formula belongs to $\mathcal{S}$.
Subcase 2.2: Suppose that $s\leq 3^{n}$. As in the previous subcase, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{s}\leftrightarrow\theta$ where $\theta$ is the consequent of $A^{i}_{s}$.
Now $A^{i}$ itself is a conjunction of conditionals which is defined to include $A^{i}_{s}$ if and only if $s\in\mathcal{S}^{\Phi}$. Let’s split into subsubcases depending on whether $s\in\mathcal{S}^{\Phi}$.
Subsubcase 2.2.1: Suppose $s\in\mathcal{S}^{\Phi}$. Then $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{s}$, whence $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta$. We reason as follows:
$\varphi_{i}\text{ is }T(x_{i})$ $\Longleftrightarrow\mathbf{FIX},w\Vdash_{V^{\star}}T(x_{i})$
$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}\mathsf{True}(\ulcorner A^{i}\urcorner)$
$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}A^{i}$
$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}\theta$
$\Longleftrightarrow\theta\text{ is }0=0$
$\Longleftrightarrow\Phi(s)_{i}=T(x_{i})$
Exactly analogous arguments deliver exactly analogous conclusions. Summarizing, we have:
$\varphi_{i}\text{ is }T(x_{i})$ $\Longleftrightarrow\Phi(s)_{i}=T(x_{i})$
$\varphi_{i}\text{ is }F(x_{i})$ $\Longleftrightarrow\Phi(s)_{i}=F(x_{i})$
$\varphi_{i}\text{ is }N(x_{i})$ $\Longleftrightarrow\Phi(s)_{i}=N(x_{i})$
These equivalences show that $\varphi=\Phi(s)$. It is the assumption of Subsubcase 2.2.1 that $s\in\mathcal{S}^{\Phi}$, i.e., $\Phi(s)\in\mathcal{S}$, so $\varphi\in\mathcal{S}$, which is desired result.
Subsubcase 2.2.2: Suppose $s\notin\mathcal{S}^{\Phi}$. Then $A^{i}_{s}$ is $C_{s}\to\lambda$. So $\mathbb{N},w\uparrow_{\mathsf{K3}}A^{i}_{s}$, whence $\mathbb{N},w\uparrow_{\mathsf{K3}}A^{i}$. As before, we conclude that $\mathbf{FIX},w\Vdash_{V^{\star}}N(x_{i})$. So $\varphi$ is $N(x_{1})\wedge\dots\wedge N(x_{n})$. By the assumption of Case (II), $\varphi\in\mathcal{S}$.
This completes the proof of Claim 3. ∎
Claim 2 and Claim 3 jointly yield Case (II) of (1) implies (2). This completes the proof of Lemma 5.9. ∎
5.5. Completeness
Combining Lemma 5.9 (the analysis of $\mathbf{FIX}$-satisfiability in terms of the prime conditions) with Carnap’s normal form theorem (Theorem 4.9), yields a complete axiomatization of the modal theory of $\mathbf{FIX}$.
Theorem 5.10.
For all $n$-formulas $\varphi\in\mathcal{L}_{\Box}$, the following are equivalent:
  1. (i)
    $\mathbf{FIX}\Vdash\varphi$
  2. (ii)
    $\mathbf{S5[Con,Ground,Min_{n}]}\vdash\varphi$
Proof.
The soundness direction is trivial. Let’s focus on completeness. Suppose that $\mathbf{S5[Con,Ground,Min_{n}]}\nvdash\varphi$. By the normal form theorem,
$\mathbf{S5[Con,Ground,Min_{n}]}\vdash\neg\varphi\leftrightarrow(\psi_{1}\vee\dots\psi_{k})$
where each $\psi_{i}$ is a disjunction of $n$-isolators. Since $\psi$ is consistent, at least one of the disjuncts $\psi_{i}$ is also consistent. It follows that $\psi_{i}$ meets the prime conditions. By Lemma 5.9, there is an $\mathcal{L}_{T}$ realization $\star$ such that
$\mathbf{FIX}\Vdash_{V^{\star}}\psi_{i}.$
Hence, we infer that $\mathbf{FIX}\nVdash\varphi.$
5.6. Counting Definable Relations
Lemma 5.9 yields an explicit characterization of the $\mathbf{FIX}$-definable binary relations, namely, each $\mathbf{FIX}$-definable binary relation is a disjunction of intensional 2-isolators. The following counting result follows from this characterization.
Theorem 5.11.
The number of $\mathbf{FIX}$-definable binary relations is exactly $2^{276}$.
Proof.
First we need to determine the number of intensional 2-isolators. By Lemma 5.9, this is to count the number of subsets of the $3\times 3$ matrix meeting the prime conditions.
There are $2^{3^{2}}$ subsets of the $3\times 3$ matrix in total. We must subtract those sets that do not meet the prime conditions:
  1. (1)
    The empty set. Subtract 1.
  2. (2)
    The number of sets that intersect the first two rows but not the third is $(2^{3}-1)^{2}=49$. Indeed, we must pick at least one element from the first row and at least one from the second and none from the third. The first row has three cells. Hence it has $72^{3}-1=$ 7 non-empty subsets. Likewise for the second. So we must subtract $497^{2}=$ 49 total.
  3. (3)
    The number of sets that intersect the first two columns but not the third is likewise $(2^{3}-1)^{2}=49$.
  4. (4)
    We must add back in all those sets that violated both the rows condition and the columns condition. Note that if a set intersects the first two rows but not the third and the first two columns but not the third then it is a subset of the upper left $2\times 2$ square. Exactly $7$ of these sets intersects the first two rows and first two columns.
  5. (5)
    The sets that intersect row 3 and column 3 but not (3,3). There are $32^{2}-1=$ 3 non-empty subsets of the first two slots of the 3rd row. Multiply this by the $3$ ways of doing the same for the 3rd column. Then multiply this with the number of ways of filling in the upper left $2\times 2$ square, which is $162^{4}=$ 16. This makes for $3\times 3\times 16$ total.
Here is the calculation:
$2762^{9}-1-49\times 2+7-3\times 3\times 16=$ 276
So there are $276$ intensional 2-isolators. The definable binary relations are exactly the disjunctions of intensional 2-isolators, so there are $2^{276}$ of them. This completes the proof. ∎
How many $\mathbf{FIX}$-definable $n$-ary relations are there? One could presumably concoct a formula using the inclusion–exclusion principle. Yet it would probably be difficult to write down an informative one. The following coarse estimate proposition follows immediately:
Proposition 5.12.
The number of $\mathbf{FIX}$-definable $n$-ary relations is $2^{k}$ for some $k$ in the interval $(2^{3^{n}-1},2^{{3^{n}}})$.
Proof.
A definable $n$-ary relation is a disjunction of intensional $n$-isolators. So this number is $2^{k}$ where $k$ is the disjunction of intensional $n$-isolators. So now we will calculate bounds on $k$.
Each intensional $n$-isolator is a subset of the set of extensional $n$-isolators, i.e., subset of the $3^{n}$ tensor. This yields the upper bound $2^{3^{n}}$.
Every subset of the $3^{n}$ tensor that includes $(3,\dots,3)$ is an intensional $n$-isolator. There are $2^{3^{n}-1}$ such subsets. This yields the lower bound. ∎
Clearly this interval is huge (and only grows as $n$ grows). One could perhaps use the inclusion–exclusion principle to count exactly the number of $\mathbf{FIX}$-definable $n$-ary relations, but it is not clear whether the resulting formula would be at all informative. Instead of focusing on the exact number, one could aim to find a more informative estimate, but we will not pursue this further in this paper.
6. Conclusions
The results in this paper concern extensions of the modal logic $\mathbf{S5}$. Many of the interesting phenomena that arise in modal logic do not arise in $\mathbf{S5}$. $\mathbf{S5}$ has Kripke semantics, but the accessibility relation is not so interesting: Often we think of the worlds as just forming an unstructured set (as we do in this paper). Moreover, as [carnap1946modalities] observed, neighboring modalities collapse into one modal, so there is no non-trivial iteration of modality in $\mathbf{S5}$. So a pessimist might conclude that $\mathbf{S5}$ is somewhat boring. The results presented in this paper suggest a different perspective, namely, that by precluding some interesting phenomena, $\mathbf{S5}$ makes room for others.
In support of this point, let me sketch an analogy with first-order model theory. Carnap’s normal form theorem resembles some classic results of first-order model theory. Indeed, normal form theorems and attendant characterizations of definable sets are standard goals in model theory when they are possible. Recall that model theorists often divide first-order theories into the tame and the wild. The following contrasting theorems give some sense of the divide between the tame and the wild:
Theorem 6.1 ([godel1931formal]).
The first-order theory of $\mathbb{Z}$ as a ring is not decidable.
Theorem 6.2 ([robinson1949definability]).
The first-order theory of $\mathbb{Q}$ as a field is not decidable.
Theorem 6.3 (Tarski 1930).
The first-order theory of $\mathbb{R}$ as a field is decidable. [^2] See [tarski1998decision].
The decidable/undecidable dichotomy is not exactly the same as the model-theoretic tame/wild dichotomy, but in these particular instances they coincide. In particular, Tarski’s theorem follows from o-minimality. A structure is o-minimal if every definable set is a finite union of intervals and a theory is o-minimal if all its structures are. Of course, there is much that o-minimal theories cannot express. The first-order theory of the reals, for instance, cannot define the set of integers or define the sine function. However, this permits logicians to give sharp, detailed characterizations of definable sets.
[…]
Sometimes wild theories like arithmetic have fragments that are both tame and rich. One method for isolating tame fragments of wild theories is pruning the signature (i.e., prohibiting some symbols from appearing in the formulas). A famous instance of this is Presburger’s theorem (contrast with Gödel’s theorem):
Theorem 6.4 ([presburger1929uber]).
The theory of $\mathbb{Z}$ as an additive group is decidable.
Presburger’s theory is tame according to various model-theoretic criteria, e.g., it eliminates quantifiers.
Another method for isolating tame fragments of wild theories is to axiomatize the behavior of certain formulas using modal logic. Provability logic is perhaps the most famous instance of this phenomenon. First-order $\mathbf{PA}$ is a wild system. Much research in the meta-theory of $\mathbf{PA}$ concerns the behavior of the provability predicate $Prov_{\mathsf{PA}}(\cdot)$ within $\mathbf{PA}$. Solovay showed that this fragment of $\mathbf{PA}$ is tame, in the following sense:
Theorem 6.5 ([solovay1976provability]).
The behavior of the predicate $Prov_{\mathsf{PA}}(\cdot)$ within $\mathbf{PA}$ is axiomatized by the decidable modal logic $\mathbf{GL}$.
The present project is another project in this spirit, although with a focus on definability rather than axiomatization. Kripke’s work of truth is straightforwardly formalizable in second-order arithmetic $\mathbf{PA_{2}}$ or in set theory $\mathbf{ZFC}$, but these theories are firmly on the wild side of the tame/wild divide. By isolating a modal fragment we are able to classify definable sets in a manner analogous to o-minimality (every definable set is a finite union of sets of a very special form). As a consequence we see that—though $\mathbf{S5}$ precludes some interesting phenomena—it permits very sharp results.
[…]

### v2

@@
 We characterize the modally definable relations and completely axiomatize the modal semantics.
-Thanks to Dan Appel, Reid Dale, Noah Schweber, Ted Slaman, and Kentarô Yamamoto for helpful discussion. Thanks to Neil Barton, Wes Holliday, and Andrew Lee for comments on drafts.
+Thanks to Dan Appel, Reid Dale, Noah Schweber, Ted Slaman, and Kentarô Yamamoto for helpful discussion. Thanks to Neil Barton, Wes Holliday, Andrew Lee, and Matt Mandelkern for comments on drafts.
 1. Introduction
@@
 This completeness theorem falls out of an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$.
-From this characterization, we infer that the set of intrinsic sentences is not modally indefinable.
+From this characterization, we infer that the set of intrinsic sentences is not modally definable.
 Thus, it is not a mere coincidence that Kripke does not lapse into modal language when defining intrinsicality.
@@
 In each case, consider the following question: Is there a variable assignment $\star$ such that $x_{1}^{\star}$ and $x_{2}^{\star}$ can attain all and only these combinations of values?
-Put another way, given a subset $\mathcal{S}$ of the $3\times 3$ matrix, is there a variable assignment $\star$ such that $\mathbf{FIX}_{V^{\star}}\Vdash\bigwedge\{\lozenge\varphi\mid\varphi\in\mathcal{S}\}\wedge\bigwedge\{\neg\lozenge\varphi\mid\varphi\notin\mathcal{S}\}$?
+Put another way, given a subset $\mathcal{S}$ of the $3\times 3$ matrix, is there a variable assignment $\star$ such that $\mathbf{FIX}\Vdash_{V^{\star}}\bigwedge\{\lozenge\varphi\mid\varphi\in\mathcal{S}\}\wedge\bigwedge\{\neg\lozenge\varphi\mid\varphi\notin\mathcal{S}\}$?
 TT TF TN
@@
 Proof.
-We prove the contrapositive.
-Suppose $\mathbf{FIX}\Vdash_{V^{\star}}\lozenge\varphi$.
-Then there is some $w$ such that $\mathbf{FIX},w\Vdash_{V^{\star}}\varphi$.
+We prove the contrapositive. Suppose $\mathbf{FIX}\Vdash_{V^{\star}}\lozenge\varphi$. Then there is some $w$ such that
+$(\triangle)\quad\mathbf{FIX},w\Vdash_{V^{\star}}\varphi.$
 We now break into cases depending on whether any of the $C_{s}$ is true at $w$.
@@
 Let’s split into subsubcases depending on whether $s\in\mathcal{S}^{\Phi}$.
-Subsubcase 2.2.1: Suppose $s\in\mathcal{S}^{\Phi}$.
-Then $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{s}$, whence $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta$.
+Subsubcase 2.2.1: Suppose $s\in\mathcal{S}^{\Phi}$. Then $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{s}$, where $A^{i}_{s}$ is $C_{s}\to\theta$.
+Note that, by the definition of $A^{i}_{s}$:
+$(\circledast)\quad\theta\text{ is }0=0\text{ just in case }\Phi(s)_{i}=T(x_{i}).$
+Since $\mathbb{N},w\vDash_{\mathsf{K3}}C_{s}$, we infer that:
+$(\boxminus)\quad\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta.$
 We reason as follows:
-$\varphi_{i}\text{ is }T(x_{i})$ $\Longleftrightarrow\mathbf{FIX},w\Vdash_{V^{\star}}T(x_{i})$
-$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}\mathsf{True}(\ulcorner A^{i}\urcorner)$
-$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}A^{i}$
-$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}\theta$
-$\Longleftrightarrow\theta\text{ is }0=0$
-$\Longleftrightarrow\Phi(s)_{i}=T(x_{i})$
+$\varphi_{i}\text{ is }T(x_{i})$ $\Longleftrightarrow\mathbf{FIX},w\Vdash_{V^{\star}}T(x_{i})\text{by ($\triangle)$}$
+$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}A^{i}\text{ since $x_{i}^{\star}=A^{i}$}$
+$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}\theta\text{ by ($\boxminus$)}$
+$\Longleftrightarrow\theta\text{ is }0=0\text{ since $\theta\in\{0=0,0=1,\lambda\}$}$
+$\Longleftrightarrow\Phi(s)_{i}=T(x_{i})\text{ by ($\circledast$)}$
 Exactly analogous arguments deliver exactly analogous conclusions. Summarizing, we have:
@@
 The soundness direction is trivial. Let’s focus on completeness. Suppose that $\mathbf{S5[Con,Ground,Min_{n}]}\nvdash\varphi$. By the normal form theorem,
-$\mathbf{S5[Con,Ground,Min_{n}]}\vdash\neg\varphi\leftrightarrow(\psi_{1}\vee\dots\psi_{k})$
+$\mathbf{S5[Con,Ground,Min_{n}]}\vdash\neg\varphi\leftrightarrow(\psi_{1}\vee\dots\vee\psi_{k})$
 where each $\psi_{i}$ is a disjunction of $n$-isolators.
-Since $\psi$ is consistent, at least one of the disjuncts $\psi_{i}$ is also consistent.
+Since $\neg\varphi$ is consistent, at least one of the disjuncts $\psi_{i}$ is also consistent.
 It follows that $\psi_{i}$ meets the prime conditions.
@@
 Yet it would probably be difficult to write down an informative one.
-The following coarse estimate proposition follows immediately:
+The following coarse estimate follows immediately:
 Proposition 5.12.
@@
 In particular, Tarski’s theorem follows from o-minimality.
-A structure is o-minimal if every definable set is a finite union of intervals and a theory is o-minimal if all its structures are.
-Of course, there is much that o-minimal theories cannot express.
-The first-order theory of the reals, for instance, cannot define the set of integers or define the sine function.
+A structure is o-minimal if every definable set is a finite union of intervals333‘Interval’ is understood in a broad sense to include rays and points.
+and a theory is o-minimal if all its structures are.
+Of course, there is much that cannot be expressed over o-minimal structures.
+The first-order language of rings, for instance, cannot define the set of integers or the sine function over the reals.
 However, this permits logicians to give sharp, detailed characterizations of definable sets.

### v3

@@
 bibliography.bib
-Abstract.
-In Outline of a Theory of Truth, Kripke introduces some of the central concepts of the logical study of truth and paradox.
+Abstract
+In Outline of a Theory of Truth, Kripke introduces many of the central concepts of the logical study of truth and paradox.
 He informally defines some of these—such as groundedness and paradoxicality—using modal locutions.
@@
 Though groundedness and paradoxicality are expressible in the modal language, we prove that intrinsicality—which Kripke emphasizes but does not define modally—is not.
-We characterize the modally definable relations and completely axiomatize the modal semantics.
-Thanks to Dan Appel, Reid Dale, Noah Schweber, Ted Slaman, and Kentarô Yamamoto for helpful discussion. Thanks to Neil Barton, Wes Holliday, Andrew Lee, and Matt Mandelkern for comments on drafts.
-1. Introduction
+This follows from a characterization of the modally definable sets and relations and an attendant axiomatization of the modal semantics.
+1 Introduction
 In Outline of a Theory of Truth, Kripke writes that
@@
 Is it a mere coincidence that modal locutions appear in the description of the grounded and paradoxical sentences but not of the intrinsic sentences?
-Or does it reflect some inherent discrepancy between the objects being defined?
-The standard way of studying these objects obviates this question.
+Might it instead reflect some inherent discrepancy between the sets being defined?
+The standard way of studying the relevant sets of sentences obviates this question.
 In particular, Kripke replaces the informal modal glosses on paradoxicality and groundedness with official extensional mathematical definitions.
@@
 For the purposes of taking these modal definitions “at face value,” we introduce a modal language and an attendant interpretation $\mathbf{FIX}$.
-Our modal language is propositional, but it contains structured atomic formulas (like in first-order logic) as opposed to propositional variables.
+The modal language is propositional, but it contains structured atomic formulas (like in first-order logic) as opposed to propositional variables.
 In particular, there are atomic formulas of the form:
@@
 $F(x)\quad\text{for ``$x$ is false''.}$”.
-We also use $N(x)$ is an abbreviation for $\neg T(x)\wedge\neg F(x)$. Kripke’s informal modal definitions of groundedness and paradoxicality are straightforwardly formalizable in this setting.
-We provide a complete and decidable axiomatization of the modal semantics. We introduce the following axioms:
+Kripke’s informal modal definitions of groundedness and paradoxicality are straightforwardly formalizable in this setting.
+The aforementioned interpretation $\mathbf{FIX}$ of the modal language interprets $\Box$ via unrestricted universal quantification over a set, [^2] In particular, $\Box$ is interpreted as a universal quantifier over the set of fixed points of the Kripke jump.
+By ‘unrestricted’ I just mean that the range of the quantifier is not constrained by an accessibility relation.
+whence it satisfies all principles of the normal modal logic $\mathbf{S5}$.
+In fact, $\mathbf{FIX}$ satisfies various additional principles (note that $N(x)$ is an abbreviation for $\neg T(x)\wedge\neg F(x)$):
 $\mathbf{Con}$$\quad\neg\big{(}T(x)\wedge F(x)\big{)}$
@@
 $\mathbf{Min_{n}}$ $\quad\big{(}\lozenge N(x_{1})\wedge\dots\wedge\lozenge N(x_{n})\big{)}\to\lozenge\big{(}N(x_{1})\wedge\dots\wedge N(x_{n})\big{)}$
-and define $\mathbf{S5[Con,Ground,Min_{n}]}$ as the normal modal logic that results from adjoining these axioms to $\mathbf{S5}$. Our main completeness theorem is the following:
+Let $\mathbf{S5[Con,Ground,Min_{n}]}$ be the normal modal logic that results from adjoining these axioms to $\mathbf{S5}$.
+The following fine-grained soundness and completeness theorem characterizes the modal theory of $\mathbf{FIX}$:
 Theorem.
@@
     $\mathbf{S5[Con,Ground,Min_{n}]}\vdash\varphi$
-This completeness theorem falls out of an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$.
-From this characterization, we infer that the set of intrinsic sentences is not modally definable.
+This completeness theorem engenders an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$.
+From this characterization, we extract the following answer to our motivating question:
+Theorem.
+The set of intrinsic sentences is not modally definable.
 Thus, it is not a mere coincidence that Kripke does not lapse into modal language when defining intrinsicality.
-Indeed, we are able to show that the intrinsic sentences are indiscernible within the modal language from another set, namely, the set of inevitable sentences, which we define later in the paper.
-The characterization of the modally definable sets builds on a classic normal form theorem for $\mathbf{S5}$ from [carnap1946modalities].
-According to Carnap’s theorem, every formula is $\mathbf{S5}$-equivalent to a disjunction of formulas in a very special form.
-This normal form result and its corollaries adumbrated here resemble classic results in model theory (particular, results in o-minimality and quantifier elimination).
-We conclude with a discussion of this analogy.
-2. Kripke’s Theory of Truth
+Here is the plan for the rest of the paper.
+In §2 we review the relevant features of Kripke’s theory of truth.
+In §3 we introduce the modal language, semantics, and axiom systems that will concern us throughout the paper.
+In §4 we characterize the sets that are definable using modal formulas with a single variable.
+In §5 we characterize the relations that are definable using modal formulas with multiple variable.
+2 Kripke’s Theory of Truth
 […]
@@
 As usual, $\ulcorner\varphi\urcorner$ is the Gödel number of the $\mathcal{L}_{T}$ sentence $\varphi$ according to some standard Gödel numbering.
-2.1. Strong Kleene logic
+2.1 Strong Kleene Logic
 Kripke defines his interpretations of $\mathcal{L}_{T}$ using the trivalent logic $\mathsf{K3}$, also known as the strong Kleene logic.
@@
   • $w\uparrow_{\mathsf{K3}}\forall x\;\varphi(x)$ otherwise.
-2.2. Classification of sentences
+2.2 Classification of Sentences
 Kripke introduces a large collection of 3-valued (true, false, neither) models for $\mathcal{L}_{T}$.
@@
 Each fixed point $w$ has the following properties:
-  1. (1)
+  1. 1.
     For each $\varphi$, $\varphi$ and $\mathsf{True}(\ulcorner\varphi\urcorner)$ have the same semantic value in $w$.
-  2. (2)
+  2. 2.
     $w$ is consistent, i.e., there is no $\varphi$ such that $w\Vdash_{\mathsf{K3}}\varphi$ and $w\Vdash_{\mathsf{K3}}\neg\varphi$.
-  3. (3)
+  3. 3.
     Hence, there is no $\varphi$ such that $w\Vdash_{\mathsf{K3}}\mathsf{True}(\ulcorner\varphi\urcorner)\wedge\mathsf{True}(\ulcorner\neg\varphi\urcorner)$
@@
 For instance, there are $\mathcal{L}_{T}$ sentences $\lambda,\tau,\gamma$ such that:
-  • $\lambda$ has the same semantic value as $\neg\mathsf{True}(\ulcorner\lambda\urcorner)$ in every fixed point.
-  • $\tau$ has the same semantic value as $\mathsf{True}(\ulcorner\tau\urcorner)$ in every fixed point.
-  • $\gamma$ has the same semantic value as $\mathsf{True}(\ulcorner\gamma\urcorner)\vee\neg\mathsf{True}(\ulcorner\gamma\urcorner)$ in every fixed point.
+  • $\lambda$ says $\neg\mathsf{True}(\ulcorner\lambda\urcorner)$.
+  • $\tau$ says $\mathsf{True}(\ulcorner\tau\urcorner)$.
+  • $\gamma$ says $\mathsf{True}(\ulcorner\gamma\urcorner)\vee\neg\mathsf{True}(\ulcorner\gamma\urcorner)$.
 In fact, there are many such sentences. We will sometimes appeal to this fact, which is well-known in the theory of truth literature:
@@
 The following claims follow:
-  1. (1)
+  1. 1.
     If a sentence has value true in one fixed point and false in another, then it has value neither in the minimum fixed point.
-  2. (2)
+  2. 2.
     If $\varphi_{1},\dots,\varphi_{n}$ are ungrounded, then they all have value neither in the minimum fixed point.
@@
 So $\tau\vee\neg\tau$ can only be made true by making a binary truth-value assignment that is incompatible with other possible binary truth-value assignments.
-3. Modal Systems
+3 Modal Systems
 In this section we introduce our modal framework, including the language, semantics, and axiom systems. Let’s start with a presentation of the modal language. In what follows:
@@
 By Remark 2.1 and Remark 2.4. ∎
-4. Single Variable Definability
+4 Single Variable Definability
 The motivations discussed in the introduction concerned definability of sets, such as the set of grounded sentences.
@@
 One conspicuous absence from the list is intrinsicality. This is not an oversight. One goal of this section is to show that the intrinsic sentences have no modal definition.
-4.1. Normal Forms
+4.1 Normal Forms
 One of our main tools will be a normal form theorem due to [carnap1946modalities]. Let’s begin by drawing a distinction between two kinds of formulas.
@@
 Let $\Gamma$ be a set of formulas and $T$ a modal system. A formula $\varphi$ is $\Gamma$-maximal for $T$ if each of the following holds:
-  1. (1)
+  1. 1.
     $T+\varphi$ is consistent
-  2. (2)
+  2. 2.
     $\varphi\in\Gamma$
-  3. (3)
+  3. 3.
     For every $\psi\in\Gamma$, one of the following holds:
@@
 For $T$ extending $\mathbf{S5}$:
-  1. (1)
+  1. 1.
     Every extensional 1-formula is $T$ equivalent to a disjunction of extensional 1-isolators for $T$.
-  2. (2)
+  2. 2.
     Every intensional 1-formula is $T$ equivalent to a disjunction of intensional 1-isolators for $T$.
-  3. (3)
+  3. 3.
     Every 1-formula is $T$ equivalent to a disjunction of 1-isolators for $T$.
-4.2. Counting
+4.2 Counting
 [carnap1946modalities] calculates the number of formulas up to provable equivalence in the modal logic $\mathbf{S5}$.
@@
 Let’s introduce some abbreviations:
-$B(x)$:
+$B(x)$
 $T(x)\wedge F(x)$
-$T_{only}(x)$:
+$T_{only}(x)$
 $T(x)\wedge\neg F(x)$
-$F_{only}(x)$:
+$F_{only}(x)$
 $\neg T(x)\wedge F(x)$
-$N(x)$:
+$N(x)$
 $\neg T(x)\wedge\neg F(x)$
@@
 ∎
-4.3. Completeness
+4.3 Completeness
 In this subsection we present a complete axiomatization of the one-variable fragment of the modal theory of $\mathbf{FIX}$.
@@
 There are exactly $6$ intensional 1-isolators for $\mathbf{S5[Con,Ground]}$. $642^{6}=$ 64. The result then follows from Proposition 4.15. ∎
-4.4. Descriptive complexity
+4.4 Descriptive Complexity
 In this subsection we will finally fulfill our promise to prove the following result:
@@
 No $\Sigma^{1}_{1}$-complete in a $\Pi^{1}_{1}$ parameter set admits a $\mathsf{Boolean}\text{-}\Sigma^{1}_{1}$ definition. ∎
-4.5. Indiscernibility
+4.5 Indiscernibility
 The previous result demonstrates the non-definability of a great number of sets, but it is a coarse tool.
@@
 $\mathcal{A}$ and $\mathcal{B}$ are $\mathbf{FIX}$-indiscernible if for every $\varphi(x)\in\mathcal{L}_{\Box}$, TFAE:
-  1. (1)
+  1. 1.
     For all $\star$ with $x^{\star}\in\mathcal{A}$, $\mathbf{FIX}\Vdash_{\mathbf{V^{\star}}}\varphi(x)$.
-  2. (2)
+  2. 2.
     For all $\star$ with $x^{\star}\in\mathcal{B}$, $\mathbf{FIX}\Vdash_{\mathbf{V^{\star}}}\varphi(x)$.
@@
   1. (i)
-    $\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
+    $\psi_{1}(x):=\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
   2. (ii)
-    $\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
+    $\psi_{2}(x):=\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
   3. (iii)
-    $\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
+    $\psi_{3}(x):=\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
   4. (iv)
-    $\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\neg\lozenge N(x)$
+    $\psi_{4}(x):=\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\neg\lozenge N(x)$
   5. (v)
-    $\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\neg\lozenge N(x)$
+    $\psi_{5}(x):=\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\neg\lozenge N(x)$
   6. (vi)
-    $\neg\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
-Note that the set of inevitable sentences is defined by the disjunction $(ii)\vee(iii)\vee(iv)\vee(v)$.
+    $\psi_{6}(x):=\neg\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
+Note that the set of inevitable sentences is defined by the disjunction $\psi_{2}(x)\vee\psi_{3}(x)\vee\psi_{4}(x)\vee\psi_{5}(x)$.
 Suppose, for the sake of contradiction, that (2) fails.
 Then there is some $\star$ such that $y^{\star}=\psi$ is inevitable and that $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(y)$.
-So $\varphi$ must be missing one of the disjuncts required to define the inevitable sentences, i.e., (ii)–(iv).
+So $\varphi$ must be missing one of the disjuncts required to define the inevitable sentences, i.e., $\psi_{2}(x)$,…,$\psi_{4}(x)$.
 Let’s break into cases.
-Case 1: $\varphi$ is missing disjunct (ii). But $\gamma$ is intrinsic and $ii(\gamma)$. So if $x^{\star}=\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
-Case 2: $\varphi$ is missing disjunct (iii).
-But $\neg\gamma$ is intrinsic and $iii(\neg\gamma)$.
-So if $x^{\star}=\neg\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
-Case 3: $\varphi$ is missing disjunct (iv). But $0=0$ is intrinsic and $iv(0=0)$. So if $x^{\star}=0=0$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
-Case 4: $\varphi$ is missing disjunct (v). But $0=1$ is intrinsic and $v(0=1)$. So if $x^{\star}=0=1$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$. ∎
+Case 1: $\varphi$ is missing disjunct $\psi_{2}$.
+But $\gamma$ is intrinsic and $\psi_{2}(\gamma)$.
+So if $x^{\star}$ is $\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
+Case 2: $\varphi$ is missing disjunct $\psi_{3}$.
+But $\neg\gamma$ is intrinsic and $\psi_{3}(\neg\gamma)$.
+So if $x^{\star}$ is $\neg\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
+Case 3: $\varphi$ is missing disjunct $\psi_{4}$. But $0=0$ is intrinsic and $\psi_{4}(0=0)$. So if $x^{\star}$ is $0=0$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
+Case 4: $\varphi$ is missing disjunct $\psi_{5}$. But $0=1$ is intrinsic and $\psi_{5}(0=1)$. So if $x^{\star}$ is $0=1$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$. ∎
 Remark 4.25.
@@
 Hence Theorem 4.24 entails that the set of intrinsic sentences is not $\mathbf{FIX}$-definable.
-5. Multivariable Definability
+5 Multivariable Definability
 The logic and combinatorics become more involved—but also more interesting—when there are many variables.
@@
 $\mathbf{FIX}\Vdash_{V^{\star}}\varphi(x_{1},\dots,x_{n})\Longleftrightarrow\langle x_{1}^{\star},\dots,x_{n}^{\star}\rangle\in\mathcal{A}.$
-5.1. Counting Formulas
+5.1 Counting Formulas
 Let’s begin with some relatively easy counting problems. We will turn to more difficult counting problems in a later subsection.
-5.1.1. The Bare Modal System
+5.1.1 The Bare Modal System
 Carnap’s normal form theorems generalize to $n$ variables. We summarize the important information for $\mathbf{S5}$ with the following bullets:
@@
 For $T$ extending $\mathbf{S5}$:
-  1. (1)
+  1. 1.
     Every extensional $n$-formula is $T$ equivalent to a disjunction of extensional $n$-isolators for $T$.
-  2. (2)
+  2. 2.
     Every intensional $n$-formula is $T$ equivalent to a disjunction of intensional $n$-isolators for $T$.
-  3. (3)
+  3. 3.
     Every $n$-formula is $T$ equivalent to a disjunction of $n$-isolators for $T$.
@@
 So there are $4^{n}\times 2^{(4^{n}-1)}$ many $n$-isolators. So the number of formulas up to provable equivalence is $2^{(4^{n}2^{(4^{n}-1)})}$. ∎
-5.1.2. With the Consistency Axiom
+5.1.2 With the Consistency Axiom
 Let’s turn out attention to $\mathbf{S5[Con]}$. We summarize the important information in the following bullets:
@@
 So there are $3^{n}\times 2^{(3^{n}-1)}$ many $n$-isolators. So the number of formulas up to provable equivalence is $2^{(3^{n}2^{(3^{n}-1)})}$. ∎
-5.2. Matrices
+5.2 Matrices
 The combinatorics of definability is much harder when $n>1$. In this section, for the sake of developing intuitions, we focus on the 2-variable case. Let’s begin by considering this matrix:
@@
 NT NF NN
-Table 1. All extensional 2-isolators.
+Table 1: All extensional 2-isolators.
 We identify each cell of the matrix with an extensional 2-isolator.
@@
 3,1 3,2 3,3
-Table 2. Standard numbering of the $3\times 3$ matrix.
+Table 2: Standard numbering of the $3\times 3$ matrix.
 Each intensional 2-isolator for $\mathbf{S5[Con]}$ corresponds to a subset of the $3\times 3$ matrix. In particular, each intensional 2-isolator has the form:
@@
 TT TF TN
-Table 3. A proper subset of the extensional 2-isolators.
+Table 3: A proper subset of the extensional 2-isolators.
 TT TF
-Table 4. A proper subset of the extensional 2-isolators.
+Table 4: A proper subset of the extensional 2-isolators.
 TT
@@
 NN
-Table 5. A proper subset of the extensional 2-isolators.
+Table 5: A proper subset of the extensional 2-isolators.
 TT TN
@@
 NT
-Table 6. A proper subset of the extensional 2-isolators.
+Table 6: A proper subset of the extensional 2-isolators.
 The answers are given here:
@@
 This last example identifies the only constraint on intensional 2-isolators that is not codified in the $\mathbf{S5[Con,Ground]}$ axioms.
-5.3. The Prime Conditions
+5.3 The Prime Conditions
 For the full completeness theorem we need new axioms:
@@
 A tensor $\mathcal{S}\subseteq[3]^{n}$ meets the prime conditions if:
-  1. (1)
+  1. 1.
     $\mathcal{S}$ is non-empty.
-  2. (2)
+  2. 2.
     For all $j\leq n$, if $\mathcal{S}$ intersects $j_{1}$ and $j_{2}$, then $\mathcal{S}$ intersects $j_{3}$.
-  3. (3)
+  3. 3.
     For all $l\leq n$, for any $t$ slices $s^{1},\dots,s^{t}$, if $\mathcal{S}$ intersects $s^{i}_{3}$ for each $i\leq t$, then $\mathcal{S}$ intersects $s^{1}_{3}\cap\dots\cap s^{t}_{3}$.
@@
 A set $\mathcal{S}$ of extensional 2-isolators meets the prime conditions if:
-  1. (1)
+  1. 1.
     $\mathcal{S}$ is non-empty.
-  2. (2)
+  2. 2.
     1. (a)
@@
       If $\mathcal{S}$ intersects column 1 and column 2, it intersects column 3.
-  3. (3)
+  3. 3.
     If $\mathcal{S}$ intersects row 3 and column 3, it contains (3,3).
-5.4. The Main Lemma
+5.4 The Main Lemma
 The main technical lemma of this paper connects the prime conditions with $\mathbf{FIX}$-satisfiability.
@@
 For $\mathcal{S}\subseteq[3]^{n}$, TFAE:
-  1. (1)
+  1. 1.
     $\mathcal{S}$ meets the prime conditions.
-  2. (2)
+  2. 2.
     There is an $\mathcal{L}_{T}$ realization $\star$ such that
@@
 Claim 2 and Claim 3 jointly yield Case (II) of (1) implies (2). This completes the proof of Lemma 5.9. ∎
-5.5. Completeness
+5.5 Completeness
 Combining Lemma 5.9 (the analysis of $\mathbf{FIX}$-satisfiability in terms of the prime conditions) with Carnap’s normal form theorem (Theorem 4.9), yields a complete axiomatization of the modal theory of $\mathbf{FIX}$.
@@
 Hence, we infer that $\mathbf{FIX}\nVdash\varphi.$
-5.6. Counting Definable Relations
+5.6 Counting Definable Relations
 Lemma 5.9 yields an explicit characterization of the $\mathbf{FIX}$-definable binary relations, namely, each $\mathbf{FIX}$-definable binary relation is a disjunction of intensional 2-isolators.
@@
 There are $2^{3^{2}}$ subsets of the $3\times 3$ matrix in total. We must subtract those sets that do not meet the prime conditions:
-  1. (1)
+  1. 1.
     The empty set. Subtract 1.
-  2. (2)
+  2. 2.
     The number of sets that intersect the first two rows but not the third is $(2^{3}-1)^{2}=49$.
@@
 So we must subtract $497^{2}=$ 49 total.
-  3. (3)
+  3. 3.
     The number of sets that intersect the first two columns but not the third is likewise $(2^{3}-1)^{2}=49$.
-  4. (4)
+  4. 4.
     We must add back in all those sets that violated both the rows condition and the columns condition.
@@
 Exactly $7$ of these sets intersects the first two rows and first two columns.
-  5. (5)
+  5. 5.
     The sets that intersect row 3 and column 3 but not (3,3).
@@
 Instead of focusing on the exact number, one could aim to find a more informative estimate, but we will not pursue this further in this paper.
-6. Conclusions
-The results in this paper concern extensions of the modal logic $\mathbf{S5}$.
-Many of the interesting phenomena that arise in modal logic do not arise in $\mathbf{S5}$.
-$\mathbf{S5}$ has Kripke semantics, but the accessibility relation is not so interesting: Often we think of the worlds as just forming an unstructured set (as we do in this paper).
-Moreover, as [carnap1946modalities] observed, neighboring modalities collapse into one modal, so there is no non-trivial iteration of modality in $\mathbf{S5}$.
-So a pessimist might conclude that $\mathbf{S5}$ is somewhat boring.
-The results presented in this paper suggest a different perspective, namely, that by precluding some interesting phenomena, $\mathbf{S5}$ makes room for others.
-In support of this point, let me sketch an analogy with first-order model theory.
-Carnap’s normal form theorem resembles some classic results of first-order model theory.
-Indeed, normal form theorems and attendant characterizations of definable sets are standard goals in model theory when they are possible.
-Recall that model theorists often divide first-order theories into the tame and the wild.
-The following contrasting theorems give some sense of the divide between the tame and the wild:
-Theorem 6.1 ([godel1931formal]).
-The first-order theory of $\mathbb{Z}$ as a ring is not decidable.
-Theorem 6.2 ([robinson1949definability]).
-The first-order theory of $\mathbb{Q}$ as a field is not decidable.
-Theorem 6.3 (Tarski 1930).
-The first-order theory of $\mathbb{R}$ as a field is decidable. [^2] See [tarski1998decision].
-The decidable/undecidable dichotomy is not exactly the same as the model-theoretic tame/wild dichotomy, but in these particular instances they coincide.
-In particular, Tarski’s theorem follows from o-minimality.
-A structure is o-minimal if every definable set is a finite union of intervals333‘Interval’ is understood in a broad sense to include rays and points.
-and a theory is o-minimal if all its structures are.
-Of course, there is much that cannot be expressed over o-minimal structures.
-The first-order language of rings, for instance, cannot define the set of integers or the sine function over the reals.
-However, this permits logicians to give sharp, detailed characterizations of definable sets.
 […]
-Sometimes wild theories like arithmetic have fragments that are both tame and rich.
-One method for isolating tame fragments of wild theories is pruning the signature (i.e., prohibiting some symbols from appearing in the formulas).
-A famous instance of this is Presburger’s theorem (contrast with Gödel’s theorem):
-Theorem 6.4 ([presburger1929uber]).
-The theory of $\mathbb{Z}$ as an additive group is decidable.
-Presburger’s theory is tame according to various model-theoretic criteria, e.g., it eliminates quantifiers.
-Another method for isolating tame fragments of wild theories is to axiomatize the behavior of certain formulas using modal logic.
-Provability logic is perhaps the most famous instance of this phenomenon.
-First-order $\mathbf{PA}$ is a wild system.
-Much research in the meta-theory of $\mathbf{PA}$ concerns the behavior of the provability predicate $Prov_{\mathsf{PA}}(\cdot)$ within $\mathbf{PA}$.
-Solovay showed that this fragment of $\mathbf{PA}$ is tame, in the following sense:
-Theorem 6.5 ([solovay1976provability]).
-The behavior of the predicate $Prov_{\mathsf{PA}}(\cdot)$ within $\mathbf{PA}$ is axiomatized by the decidable modal logic $\mathbf{GL}$.
-The present project is another project in this spirit, although with a focus on definability rather than axiomatization.
-Kripke’s work of truth is straightforwardly formalizable in second-order arithmetic $\mathbf{PA_{2}}$ or in set theory $\mathbf{ZFC}$, but these theories are firmly on the wild side of the tame/wild divide.
-By isolating a modal fragment we are able to classify definable sets in a manner analogous to o-minimality (every definable set is a finite union of sets of a very special form).
-As a consequence we see that—though $\mathbf{S5}$ precludes some interesting phenomena—it permits very sharp results.
-[…]

### v4

@@
 bibliography.bib
-Abstract
+Abstract.
 In Outline of a Theory of Truth, Kripke introduces many of the central concepts of the logical study of truth and paradox.
@@
 This follows from a characterization of the modally definable sets and relations and an attendant axiomatization of the modal semantics.
-1 Introduction
+Thanks to Dan Appel, Reid Dale, Noah Schweber, Ted Slaman, and Kentarô Yamamoto for helpful discussion. Thanks to Neil Barton, Wes Holliday, Andrew Lee, and Matthew Mandelkern for comments on drafts.
+1. Introduction
 In Outline of a Theory of Truth, Kripke writes that
@@
 In particular, there are atomic formulas of the form:
-$T(x)\quad\text{for ``$x$ is true''}$”
-$F(x)\quad\text{for ``$x$ is false''.}$”.
+$T(x)\quad\text{for ``$x$ is true.''}$.”
+$F(x)\quad\text{for ``$x$ is false.''}$.”
 Kripke’s informal modal definitions of groundedness and paradoxicality are straightforwardly formalizable in this setting.
+[^2] It is is worth noting that there is other work mixing modality and Kripke’s theory of truth, see [stern2014modality, nicolai2021modal, cook2022outline].
+However, none of these address the particular questions addressed here.
 The aforementioned interpretation $\mathbf{FIX}$ of the modal language interprets $\Box$ via unrestricted universal quantification over a set, [^3] In particular, $\Box$ is interpreted as a universal quantifier over the set of fixed points of the Kripke jump.
@@
 $\mathbf{Con}$$\quad\neg\big{(}T(x)\wedge F(x)\big{)}$
-$\mathbf{Ground}$ $\quad(\lozenge T(x)\wedge\lozenge F(x))\to\lozenge N(x)$
+$\mathbf{Ground}$ $\quad\big{(}\lozenge T(x)\wedge\lozenge F(x)\big{)}\to\lozenge N(x)$
 $\mathbf{Min_{n}}$ $\quad\big{(}\lozenge N(x_{1})\wedge\dots\wedge\lozenge N(x_{n})\big{)}\to\lozenge\big{(}N(x_{1})\wedge\dots\wedge N(x_{n})\big{)}$
-Let $\mathbf{S5[Con,Ground,Min_{n}]}$ be the normal modal logic that results from adjoining these axioms to $\mathbf{S5}$.
+Let $\mathbf{S5[Con,Ground,Min_{n}]}$ be the modal system that results from adjoining these axioms to $\mathbf{S5}$.
 The following fine-grained soundness and completeness theorem characterizes the modal theory of $\mathbf{FIX}$:
 Theorem.
-For all formulas $\varphi$ with $n$ variables, the following are equivalent:
+For all formulas $\varphi$ with at most $n$ variables, the following are equivalent:
   1. (i)
@@
     $\mathbf{S5[Con,Ground,Min_{n}]}\vdash\varphi$
-This completeness theorem engenders an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$.
+This completeness theorem emerges from an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$.
 From this characterization, we extract the following answer to our motivating question:
@@
 In §5 we characterize the relations that are definable using modal formulas with multiple variable.
-2 Kripke’s Theory of Truth
+2. Kripke’s Theory of Truth
 […]
@@
 As usual, $\ulcorner\varphi\urcorner$ is the Gödel number of the $\mathcal{L}_{T}$ sentence $\varphi$ according to some standard Gödel numbering.
-2.1 Strong Kleene Logic
+2.1. Strong Kleene Logic
 Kripke defines his interpretations of $\mathcal{L}_{T}$ using the trivalent logic $\mathsf{K3}$, also known as the strong Kleene logic.
@@
   • $w\uparrow_{\mathsf{K3}}\forall x\;\varphi(x)$ otherwise.
-2.2 Classification of Sentences
+2.2. Classification of Sentences
 Kripke introduces a large collection of 3-valued (true, false, neither) models for $\mathcal{L}_{T}$.
@@
 The details of this function are not important for us, but the “fixed points” terminology is ubiquitous, so we will continue to use it.
+For work on the structure of these fixed points see [cantini1989notes, fitting1986notes].
 Remark 2.1.
 Each fixed point $w$ has the following properties:
-  1. 1.
+  1. (1)
     For each $\varphi$, $\varphi$ and $\mathsf{True}(\ulcorner\varphi\urcorner)$ have the same semantic value in $w$.
-  2. 2.
-    $w$ is consistent, i.e., there is no $\varphi$ such that $w\Vdash_{\mathsf{K3}}\varphi$ and $w\Vdash_{\mathsf{K3}}\neg\varphi$.
-  3. 3.
+  2. (2)
+    $w$ is consistent, i.e., there is no $\varphi$ such that $w\vDash_{\mathsf{K3}}\varphi$ and $w\vDash_{\mathsf{K3}}\neg\varphi$ [^4] Like Kripke, we restrict our attention to consistent fixed points.
+It is now common to study all fixed points without this restriction.
+  3. (3)
     Hence, there is no $\varphi$ such that $w\vDash_{\mathsf{K3}}\mathsf{True}(\ulcorner\varphi\urcorner)\wedge\mathsf{True}(\ulcorner\neg\varphi\urcorner)$
-For any formula $\varphi(x)$ in the language of $\mathcal{L}_{T}$, we can formulate an $\mathcal{L}_{T}$ sentence $\psi$ which “says” $\varphi(\ulcorner\psi\urcorner)$.
+For any formula $\varphi(x)$ in the language of $\mathcal{L}_{T}$, a standard self-reference construction yields an $\mathcal{L}_{T}$ sentence $\psi$ which “says” $\varphi(\ulcorner\psi\urcorner)$.
 In particular, $\psi$ and $\varphi(\ulcorner\psi\urcorner)$ have the same semantic value in each fixed point.
@@
 Remark 2.2.
-There are infinitely many logically independent truth-teller sentences. That is, there is an infinite collection $T$ of sentences such that:
-  • For all $\tau\in T$, $\tau$ has the same semantic value as $\mathsf{True}(\ulcorner\tau\urcorner)$ in every fixed point.
-  • For any sequence $\tau_{1},\dots,\tau_{n}$ of sentences from $T$ and any assignment of true, false, and neither to $\tau_{1},\dots,\tau_{n}$, there is a fixed point realizing exactly that truth-value assignment.
+There are infinitely many logically independent truth-teller sentences.
+Hence, there is an infinite collection $T$ of sentences such that for any sequence $\tau_{1},\dots,\tau_{n}$ of sentences from $T$ and any assignment of true, false, and neither to $\tau_{1},\dots,\tau_{n}$, there is a fixed point realizing exactly that truth-value assignment.
 Definition 2.3.
@@
 The following claims follow:
-  1. 1.
+  1. (1)
     If a sentence has value true in one fixed point and false in another, then it has value neither in the minimum fixed point.
-  2. 2.
+  2. (2)
     If $\varphi_{1},\dots,\varphi_{n}$ are ungrounded, then they all have value neither in the minimum fixed point.
@@
 So $\tau\vee\neg\tau$ can only be made true by making a binary truth-value assignment that is incompatible with other possible binary truth-value assignments.
-3 Modal Systems
+3. Modal Systems
 In this section we introduce our modal framework, including the language, semantics, and axiom systems. Let’s start with a presentation of the modal language. In what follows:
@@
   1. (i)
-    $V^{\star}_{1}(x)$ is the set of models in which $x^{\star}$ is true.
+    $V^{\star}_{1}(x)$ is the set of worlds in which $x^{\star}$ is true.
   2. (ii)
-    $V^{\star}_{2}(x)$ is the set of models in which $x^{\star}$ is false.
+    $V^{\star}_{2}(x)$ is the set of worlds in which $x^{\star}$ is false.
 $\mathbf{FIX}\Vdash\varphi$ just in case $\mathbf{FIX}\Vdash_{V^{\star}}\varphi$ for every $\mathcal{L}_{T}$-realization $\star$.
@@
 By Remark 2.1 and Remark 2.4. ∎
-4 Single Variable Definability
+4. Single Variable Definability
 The motivations discussed in the introduction concerned definability of sets, such as the set of grounded sentences.
@@
 One conspicuous absence from the list is intrinsicality. This is not an oversight. One goal of this section is to show that the intrinsic sentences have no modal definition.
-4.1 Normal Forms
+4.1. Normal Forms
 One of our main tools will be a normal form theorem due to [carnap1946modalities]. Let’s begin by drawing a distinction between two kinds of formulas.
@@
 Let $\Gamma$ be a set of formulas and $T$ a modal system. A formula $\varphi$ is $\Gamma$-maximal for $T$ if each of the following holds:
-  1. 1.
+  1. (1)
     $T+\varphi$ is consistent
-  2. 2.
+  2. (2)
     $\varphi\in\Gamma$
-  3. 3.
+  3. (3)
     For every $\psi\in\Gamma$, one of the following holds:
@@
 For $T$ extending $\mathbf{S5}$:
-  1. 1.
+  1. (1)
     Every extensional 1-formula is $T$ equivalent to a disjunction of extensional 1-isolators for $T$.
-  2. 2.
+  2. (2)
     Every intensional 1-formula is $T$ equivalent to a disjunction of intensional 1-isolators for $T$.
-  3. 3.
+  3. (3)
     Every 1-formula is $T$ equivalent to a disjunction of 1-isolators for $T$.
-4.2 Counting
+4.2. Counting
 [carnap1946modalities] calculates the number of formulas up to provable equivalence in the modal logic $\mathbf{S5}$.
@@
 Let’s introduce some abbreviations:
-$B(x)$
-$T(x)\wedge F(x)$
-$T_{only}(x)$
-$T(x)\wedge\neg F(x)$
-$F_{only}(x)$
-$\neg T(x)\wedge F(x)$
-$N(x)$
-$\neg T(x)\wedge\neg F(x)$
-$\mathbf{S5}$ has 4 extensional 1-isolators and 15 intensional 1-isolators, namely, all formulas of the form:
+$B(x)$ $\quad T(x)\wedge F(x)$
+$T_{only}(x)$ $\quad T(x)\wedge\neg F(x)$
+$F_{only}(x)$$\quad\neg T(x)\wedge F(x)$
+$N(x)$$\quad\neg T(x)\wedge\neg F(x)$
+$\mathbf{S5}$ has 4 extensional 1-isolators and 15 intensional 1-isolators, namely, all 1-formulas of the form:
 ¬${}^{+}_{\neg}\lozenge B(x)\wedge^{+}_{\neg}\lozenge T_{only}(x)\wedge^{+}_{\neg}\lozenge F_{only}(x)\wedge^{+}_{\neg}\lozenge N(x)$
@@
 So there are $4\times 8=32$ many 1-isolators for $\mathbf{S5}$.
-Hence, the number of formulas up to provable equivalence is exactly $2962^{32}=4{,}294{,}967{,}$ 296.
+Hence, the number of 1-formulas up to provable equivalence is exactly $2962^{32}=4{,}294{,}967{,}$ 296.
 ∎
@@
 A 1-isolator is a consistent conjunction of an extensional 1-isolator and an intensional 1-isolator.
-So there are $3+3+4=10$ many 1-isolators for $\mathbf{S5[Con]}$.
+So there are $3+3+4=10$ many 1-isolators for $\mathbf{S5[Con,Ground]}$.
 Hence, the number of formulas up to provable-equivalence is exactly $10242^{10}=$ 1024.
 ∎
-4.3 Completeness
+4.3. Completeness
 In this subsection we present a complete axiomatization of the one-variable fragment of the modal theory of $\mathbf{FIX}$.
@@
 ∎
+Before continuing, we note that the normal form theorem also yields an exact characterization of the number of FIX-definable sets.
 Proposition 4.15.
@@
 There are exactly $6$ intensional 1-isolators for $\mathbf{S5[Con,Ground]}$. $642^{6}=$ 64. The result then follows from Proposition 4.15. ∎
-4.4 Descriptive Complexity
+4.4. Descriptive Complexity
 In this subsection we will finally fulfill our promise to prove the following result:
@@
 There are multiple ways of verifying this, and we will present two. One could check all 64 single variable definitions, although this would be tedious. Luckily, there are more conceptual arguments.
-Every $\mathbf{FIX}$-definable is definable by an intensional 1-formula.
+Every $\mathbf{FIX}$-definable set is definable by an intensional 1-formula.
 This is because if $\varphi$ $\mathbf{FIX}$-defines a set $\mathcal{A}$, then so does $\Box\varphi$.
@@
 No $\Sigma^{1}_{1}$-complete in a $\Pi^{1}_{1}$ parameter set admits a $\mathsf{Boolean}\text{-}\Sigma^{1}_{1}$ definition. ∎
-4.5 Indiscernibility
-The previous result demonstrates the non-definability of a great number of sets, but it is a coarse tool.
-For instance, descriptive complexity considerations do not yield the non-definability of the arithmetic truths, since they form a $\Delta^{1}_{1}$ set, but the the arithmetic truths are not definable.
-We will now prove the indefinability of the intrinsic sentences in another way.
+4.5. Indiscernibility
+The previous result demonstrates the non-definability of a great number of sets.
+Nevertheless, it is a coarse tool.
+We can prove the indefinability of the intrinsic sentences in a more informative way.
 Informally, sets $\mathcal{A}$ and $\mathcal{B}$ are $\mathbf{FIX}$-indiscernible just in case every formula that applies to the $\mathcal{A}$s applies to the $\mathcal{B}$s and vice-versa.
@@
 $\mathcal{A}$ and $\mathcal{B}$ are $\mathbf{FIX}$-indiscernible if for every $\varphi(x)\in\mathcal{L}_{\Box}$, TFAE:
-  1. 1.
+  1. (1)
     For all $\star$ with $x^{\star}\in\mathcal{A}$, $\mathbf{FIX}\Vdash_{\mathbf{V^{\star}}}\varphi(x)$.
-  2. 2.
+  2. (2)
     For all $\star$ with $x^{\star}\in\mathcal{B}$, $\mathbf{FIX}\Vdash_{\mathbf{V^{\star}}}\varphi(x)$.
@@
 Then there is some $\star$ such that $y^{\star}=\psi$ is inevitable and that $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(y)$.
-So $\varphi$ must be missing one of the disjuncts required to define the inevitable sentences, i.e., $\psi_{2}(x)$,…,$\psi_{4}(x)$.
+So $\varphi$ must be missing one of the disjuncts required to define the inevitable sentences, i.e., $\psi_{2}(x)$,…,$\psi_{5}(x)$.
 Let’s break into cases.
@@
 Hence Theorem 4.24 entails that the set of intrinsic sentences is not $\mathbf{FIX}$-definable.
-5 Multivariable Definability
+5. Multivariable Definability
 The logic and combinatorics become more involved—but also more interesting—when there are many variables.
@@
 $\mathbf{FIX}\Vdash_{V^{\star}}\varphi(x_{1},\dots,x_{n})\Longleftrightarrow\langle x_{1}^{\star},\dots,x_{n}^{\star}\rangle\in\mathcal{A}.$
-5.1 Counting Formulas
+5.1. Counting Formulas
 Let’s begin with some relatively easy counting problems. We will turn to more difficult counting problems in a later subsection.
-5.1.1 The Bare Modal System
+5.1.1. The Bare Modal System
 Carnap’s normal form theorems generalize to $n$ variables. We summarize the important information for $\mathbf{S5}$ with the following bullets:
@@
 For $T$ extending $\mathbf{S5}$:
-  1. 1.
+  1. (1)
     Every extensional $n$-formula is $T$ equivalent to a disjunction of extensional $n$-isolators for $T$.
-  2. 2.
+  2. (2)
     Every intensional $n$-formula is $T$ equivalent to a disjunction of intensional $n$-isolators for $T$.
-  3. 3.
+  3. (3)
     Every $n$-formula is $T$ equivalent to a disjunction of $n$-isolators for $T$.
@@
 So fixing a particular extensional $n$-isolator $\varepsilon$, there are $2^{(4^{n}-1)}$ many intensional $n$-isolators with $\lozenge\varepsilon$ as a conjunct.
-So there are $4^{n}\times 2^{(4^{n}-1)}$ many $n$-isolators. So the number of formulas up to provable equivalence is $2^{(4^{n}2^{(4^{n}-1)})}$. ∎
-5.1.2 With the Consistency Axiom
+So there are $4^{n}\times 2^{(4^{n}-1)}$ many $n$-isolators. So the number of $n$-formulas up to provable equivalence is $2^{(4^{n}2^{(4^{n}-1)})}$. ∎
+5.1.2. With the Consistency Axiom
 Let’s turn out attention to $\mathbf{S5[Con]}$. We summarize the important information in the following bullets:
@@
 So fixing a particular extensional $n$-isolator $\varepsilon$, there are $2^{(3^{n}-1)}$ many intensional $n$-isolators with $\lozenge\varepsilon$ as a conjunct.
-So there are $3^{n}\times 2^{(3^{n}-1)}$ many $n$-isolators. So the number of formulas up to provable equivalence is $2^{(3^{n}2^{(3^{n}-1)})}$. ∎
-5.2 Matrices
+So there are $3^{n}\times 2^{(3^{n}-1)}$ many $n$-isolators. So the number of $n$-formulas up to provable equivalence is $2^{(3^{n}2^{(3^{n}-1)})}$. ∎
+5.2. Matrices
 The combinatorics of definability is much harder when $n>1$. In this section, for the sake of developing intuitions, we focus on the 2-variable case. Let’s begin by considering this matrix:
@@
 NT NF NN
-Table 1: All extensional 2-isolators.
+Table 1. All extensional 2-isolators.
 We identify each cell of the matrix with an extensional 2-isolator.
@@
 3,1 3,2 3,3
-Table 2: Standard numbering of the $3\times 3$ matrix.
+Table 2. Standard numbering of the $3\times 3$ matrix.
 Each intensional 2-isolator for $\mathbf{S5[Con]}$ corresponds to a subset of the $3\times 3$ matrix. In particular, each intensional 2-isolator has the form:
@@
 TT TF TN
-Table 3: A proper subset of the extensional 2-isolators.
+Table 3. A proper subset of the extensional 2-isolators.
 TT TF
-Table 4: A proper subset of the extensional 2-isolators.
+Table 4. A proper subset of the extensional 2-isolators.
 TT
@@
 NN
-Table 5: A proper subset of the extensional 2-isolators.
+Table 5. A proper subset of the extensional 2-isolators.
 TT TN
@@
 NT
-Table 6: A proper subset of the extensional 2-isolators.
+Table 6. A proper subset of the extensional 2-isolators.
 The answers are given here:
@@
 This last example identifies the only constraint on intensional 2-isolators that is not codified in the $\mathbf{S5[Con,Ground]}$ axioms.
-5.3 The Prime Conditions
+5.3. The Prime Conditions
 For the full completeness theorem we need new axioms:
@@
 For instance, to see that $\mathbf{Min_{3}}\vdash\mathbf{Min_{2}}$ note that the following instance of $\mathbf{Min_{3}}$:
-$\big{(}\lozenge N(x_{1})\wedge N(x_{1})\wedge\lozenge N(x_{2})\big{)}\to\lozenge\big{(}N(x_{1})\wedge N(x_{1})\wedge N(x_{2})\big{)}$
+$\big{(}\lozenge N(x_{1})\wedge\lozenge N(x_{1})\wedge\lozenge N(x_{2})\big{)}\to\lozenge\big{(}N(x_{1})\wedge N(x_{1})\wedge N(x_{2})\big{)}$
 is (logically equivalent to) an instance of $\mathbf{Min_{2}}$.
@@
 For the $n$-variable case, we must turn from matrices to tensors, which are generalizations of matrices beyond two dimensions.
-In particular, will be interested in $3^{n}$ tensors, i.e., $\underbrace{3\times\ldots\times 3}\limits_{\mbox{\scriptsize$n$ times}}$ matrices.
-Let $[3]^{n}$ be the set of triples from the set $\{1,2,3\}$.
+In particular, we will be interested in $3^{n}$ tensors, i.e., $\underbrace{3\times\ldots\times 3}\limits_{\mbox{\scriptsize$n$ times}}$ matrices.
+Let $[3]^{n}$ be the set of $n$-tuples from the set $\{1,2,3\}$.
 This is the set of positions in a $3^{n}$ tensor.
@@
 Definition 5.6.
-For $k\leq 3$ and $j\leq n$, the $k^{th}$-layer of the $j^{th}$ slice (written $j_{k}$) is
-$\{a\in[3^{n}]\mid\exists b_{1},\dots,\exists b_{n-1}\;a=\langle b_{1},\dots b_{j-1},k,b_{j},\dots,b_{n-1}\rangle\}$
+For $k\leq 3$ and $j\leq n$, the $k^{th}$ layer of the $j^{th}$ slice (written $j_{k}$) is
+$\{a\in[3]^{n}\mid\exists b_{1},\dots,\exists b_{n-1}\;a=\langle b_{1},\dots b_{j-1},k,b_{j},\dots,b_{n-1}\rangle\}$
 Note that as $j_{k}$ is defined, it is a set of positions in the $3^{n}$ tensor.
@@
 A tensor $\mathcal{S}\subseteq[3]^{n}$ meets the prime conditions if:
-  1. 1.
+  1. (1)
     $\mathcal{S}$ is non-empty.
-  2. 2.
+  2. (2)
     For all $j\leq n$, if $\mathcal{S}$ intersects $j_{1}$ and $j_{2}$, then $\mathcal{S}$ intersects $j_{3}$.
-  3. 3.
-    For all $l\leq n$, for any $t$ slices $s^{1},\dots,s^{t}$, if $\mathcal{S}$ intersects $s^{i}_{3}$ for each $i\leq t$, then $\mathcal{S}$ intersects $s^{1}_{3}\cap\dots\cap s^{t}_{3}$.
+  3. (3)
+    For all $t\leq n$, for any $t$ slices $s^{1},\dots,s^{t}$, if $\mathcal{S}$ intersects $s^{i}_{3}$ for each $i\leq t$, then $\mathcal{S}$ intersects $s^{1}_{3}\cap\dots\cap s^{t}_{3}$.
 Remark 5.8.
@@
 A set $\mathcal{S}$ of extensional 2-isolators meets the prime conditions if:
-  1. 1.
+  1. (1)
     $\mathcal{S}$ is non-empty.
-  2. 2.
+  2. (2)
     1. (a)
@@
       If $\mathcal{S}$ intersects column 1 and column 2, it intersects column 3.
-  3. 3.
+  3. (3)
     If $\mathcal{S}$ intersects row 3 and column 3, it contains (3,3).
-5.4 The Main Lemma
+5.4. The Main Lemma
 The main technical lemma of this paper connects the prime conditions with $\mathbf{FIX}$-satisfiability.
@@
 For $\mathcal{S}\subseteq[3]^{n}$, TFAE:
-  1. 1.
+  1. (1)
     $\mathcal{S}$ meets the prime conditions.
-  2. 2.
+  2. (2)
     There is an $\mathcal{L}_{T}$ realization $\star$ such that
@@
 Note that $\varphi$ is a conjunction of the form $\varphi_{1}(x_{1})\wedge\dots\wedge\varphi_{n}(x_{n})$.
-Note also that $\Phi^{-1}(\varphi)$ is some number $k\leq 3^{n}$.
+Note also that $\Phi^{-1}(\varphi)$ is some number $k\leq 3^{n}$ (by $\Phi^{-1}$ I mean the inverse of $\Phi$; so $\Phi^{-1}(\varphi)$ is just the number that $\Phi$ maps to $\varphi$).
 Since $C_{k}$ is a Boolean statement about independent truth-tellers, there is some $w$ such that $\mathbb{N},w\vDash_{\mathsf{K3}}C_{k}$ (see Remark 2.2).
@@
 Which is to say that for all conjuncts $A^{i}_{j}$ of $A^{i}$ except for $A^{i}_{k}$, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{j}$.
-Thus, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{k}$.
-Since $\mathbb{N},w\vDash_{\mathsf{K3}}C_{k}$, we infer that, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta$.
+Thus, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{k}$.
+Since $\mathbb{N},w\vDash_{\mathsf{K3}}C_{k}$, we infer that $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}\theta$.
 We now want to prove that $\mathbf{FIX},w\Vdash_{V^{\star}}\varphi_{i}$, where $\varphi_{i}$ is the $i$th conjunct of $\varphi$. Let’s break into cases depending on which formula $\varphi_{i}$ is.
@@
 We now break into cases depending on whether any of the $C_{s}$ is true at $w$.
-Case 1: Suppose $\mathbb{N},w\uparrow_{\mathsf{K3}}C_{s}$ for all $s\leq 16$. We claim that $A^{1},\dots,A^{n}$ all have value neither in $w$.
+Case 1: Suppose $\mathbb{N},w\uparrow_{\mathsf{K3}}C_{s}$ for all $s\leq 2^{k}$. We claim that $A^{1},\dots,A^{n}$ all have value neither in $w$.
 To see this, note that $A^{i}$ is a conjunction of conditionals each of which has an antecedent $C_{s}$, which has value neither.
@@
 $A^{i}$ is a conjunction of conditionals of the form $C_{j}\to\theta^{A}_{i}$ such that for $j\neq s$, $\mathbb{N},w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}C_{j}$, whence $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{j}$.
-Accordingly, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta$, where $\theta$ is the consequent of $A^{i}_{s}$.
+Accordingly, $\mathbb{N},w\uparrow_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\uparrow_{\mathsf{K3}}\theta$, where $\theta$ is the consequent of $A^{i}_{s}$.
 Since $s>3^{n}$, $\theta$ is $\lambda$.
@@
 Thus, $\varphi$ is $N(x_{1})\wedge\dots\wedge N(x_{n})$, i.e., (3, …, 3). Once again, since we are working in Case (II), we have assumed that this formula belongs to $\mathcal{S}$.
-Subcase 2.2: Suppose that $s\leq 3^{n}$. As in the previous subcase, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{s}\leftrightarrow\theta$ where $\theta$ is the consequent of $A^{i}_{s}$.
+Subcase 2.2: Suppose that $s\leq 3^{n}$.
+As in the previous subcase, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{s}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}\theta$ where $\theta$ is the consequent of $A^{i}_{s}$.
 Now $A^{i}$ itself is a conjunction of conditionals which is defined to include $A^{i}_{s}$ if and only if $s\in\mathcal{S}^{\Phi}$.
 Let’s split into subsubcases depending on whether $s\in\mathcal{S}^{\Phi}$.
-Subsubcase 2.2.1: Suppose $s\in\mathcal{S}^{\Phi}$. Then $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{s}$, where $A^{i}_{s}$ is $C_{s}\to\theta$.
+Subsubcase 2.2.1: Suppose $s\in\mathcal{S}^{\Phi}$. Then $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{s}$, where $A^{i}_{s}$ is $C_{s}\to\theta$.
 Note that, by the definition of $A^{i}_{s}$:
@@
 Since $\mathbb{N},w\vDash_{\mathsf{K3}}C_{s}$, we infer that:
-$(\boxminus)\quad\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta.$
+$(\boxminus)\quad\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}\theta.$
 We reason as follows:
@@
 Claim 2 and Claim 3 jointly yield Case (II) of (1) implies (2). This completes the proof of Lemma 5.9. ∎
-5.5 Completeness
+5.5. Completeness
 Combining Lemma 5.9 (the analysis of $\mathbf{FIX}$-satisfiability in terms of the prime conditions) with Carnap’s normal form theorem (Theorem 4.9), yields a complete axiomatization of the modal theory of $\mathbf{FIX}$.
@@
 Hence, we infer that $\mathbf{FIX}\nVdash\varphi.$
-5.6 Counting Definable Relations
+5.6. Counting Definable Relations
 Lemma 5.9 yields an explicit characterization of the $\mathbf{FIX}$-definable binary relations, namely, each $\mathbf{FIX}$-definable binary relation is a disjunction of intensional 2-isolators.
@@
 There are $2^{3^{2}}$ subsets of the $3\times 3$ matrix in total. We must subtract those sets that do not meet the prime conditions:
-  1. 1.
+  1. (1)
     The empty set. Subtract 1.
-  2. 2.
+  2. (2)
     The number of sets that intersect the first two rows but not the third is $(2^{3}-1)^{2}=49$.
@@
 So we must subtract $497^{2}=$ 49 total.
-  3. 3.
+  3. (3)
     The number of sets that intersect the first two columns but not the third is likewise $(2^{3}-1)^{2}=49$.
-  4. 4.
+  4. (4)
     We must add back in all those sets that violated both the rows condition and the columns condition.
@@
 Exactly $7$ of these sets intersects the first two rows and first two columns.
-  5. 5.
+  5. (5)
     The sets that intersect row 3 and column 3 but not (3,3).
@@
 So there are $276$ intensional 2-isolators. The definable binary relations are exactly the disjunctions of intensional 2-isolators, so there are $2^{276}$ of them. This completes the proof. ∎
-How many $\mathbf{FIX}$-definable $n$-ary relations are there?
-One could presumably concoct a formula using the inclusion–exclusion principle.
-Yet it would probably be difficult to write down an informative one.
-The following coarse estimate follows immediately:
+How many $\mathbf{FIX}$-definable $n$-ary relations are there? The following coarse estimate is immediate:
 Proposition 5.12.